### v1

# Multistability of graphene nanobubbles

Abstract
Using the example of Ar, Kr, and Xe atoms, it is shown that graphene nanobubbles on flat substrates are multistable systems. A nanobubble can have many stable stationary states, each characterized by the number of layers, $l$, within the cluster of internal atoms. The layers are circular in shape, concentrically stacked on top of each other, forming an $l$-stepped pyramid with a flat top. The covering of this pyramid with a graphene sheet is achieved through its local stretching. The valence bonds of the sheet stretch only over the group of internal atoms; outside the coverage zone, the sheet remains undeformed and lies flush against the substrate. The maximum possible number of layers, $l_{m}$, increases monotonically with the number of atoms $N$ ($l_{m}=6$ for $N=4000$). The graphene sheet, interacting with the substrate, compresses the internal atom cluster against it, generating an internal pressure of $P\sim 1$ GPa. Numerical simulations of thermal vibrations reveal that among all $l$-layer configurations of a nanobubble, there is always one ”ground” state. Upon heating, this ground state smoothly transitions into a layerless liquid state. All other stationary states transform into this ground state once a certain temperature is reached (for $N=4000$, the ground state corresponds to state with $l=4$). The coexistence of several stable states with different numbers of layers at low temperatures leads to the absence of a universal shape for the nanobubbles. In this scenario, the height-to-radius ratio, $H/R$, can vary from 0 to 0.24, depending on the number of layers.
Keywords: Van der Walls bubble, graphene nanobubble, molecular modelling, phase transition
I Introduction
Carbon atoms can form numerous structures, among which the monoatomic crystalline layer graphene has recently attracted significant research attention Novoselov2004 ; Geim2007 ; Soldano2010 ; Baimova2014a ; Baimova2014b . This nanomaterial is of great interest due to its unique electronic Geim2009 , mechanical Lee2008 , and thermal properties Baladin2008 ; Liu2015 . Another important physical and chemical property of graphene is its ability to trap gas molecules under high pressure when placed on various flat substrates. This can lead to the formation of bubbles ranging in size from nanometers to micrometers.
High interest is also directed towards other graphene-like materials, such as sheets of hexagonal boron nitride (h-BN), fluorographene (CF), molybdenum disulfide (MoS2), and tungsten selenide (WSe2). These two-dimensional materials offer vast possibilities for creating heterostructures in the form of van der Waals stacks, which possess novel properties Geim2013 ; Wang2013 ; Xiang2020 ; Zhang2020 . An important feature of such heterostructures is their ability to trap molecules in localized regions between adjacent layers. Due to van der Waals interactions between the layers, internal pressure on the order of several GPa can arise within these traps (nanopockets) Vasu2016 ; Khestanova2016 ; Hu2023 , which can significantly change the properties of the captured material. Covering a molecular system on a flat substrate with a graphene sheet (van der Waals encapsulation) is a convenient method for creating high local pressure to modify the system’s properties Zhang2017 or for its storage Zheng2012 ; Slepchenkov2018 ; Apkadirova2022 . In graphene nanobubbles, high pressure is induced by the forces of interaction between the sheet and the substrate.
Encapsulation structures (nanobubbles) can form spontaneously as surface defects during the mechanical transfer of two-dimensional sheets in the creation of van der Waals heterostructures Khestanova2016 ; Sanchez2018 . Nanobubbles can also be intentionally fabricated using ion bombardment Blundo2020 ; Zamborlini2015 ; Villarreal2021 or selective adsorption methods Zahra2020 .
In addition to experimental methods, analytical models and numerical simulations are used to study nanobubbles. Analytical models employ elastic plate theory, taking into account the contribution of adhesion energy between the layers of van der Waals heterostructures Yue2012 ; Wang2013a ; Dai2018 . More complex models also incorporate the equation of state of the substance inside the nanobubble Zhilyaev2019 ; Iakovlev2019 ; Aslymov2020 ; Qu2023 and nonlinear effects Lyublinskaya2020 . Using such models, it has been shown that nanobubbles with a circular shape exhibit universal scaling behavior: the ratio of the nanobubble’s height to its radius, $H/R$, remains constant. Numerical simulations allow not only for the verification of analytical results but also for gaining deeper insights into nanobubble properties. It has been demonstrated that spatial confinement and internal pressure within a nanobubble promote the formation of internal crystalline structures, which can be planar in shape Iakovlev2017 , and that the nanobubble’s shape can critically depend on the properties of the encapsulated substance Ghorbanfekr2017 . Numerical simulations have shown Jain2017 that a nanobubble containing an ideal gas always has a circular shape with a constant ratio $H/R=0.204$. Molecular dynamics simulations have modeled transitions of argon inside a nanobubble into a liquid-crystal state Korneva2023 . It has been demonstrated that the presence of nanobubbles can lead to a reduction in the interfacial thermal conductivity of layered van der Waals structures Qu2023a , as well as cause their significant thermal contraction Qu2023b .
In the paper, all possible stationary states of graphene nanobubbles will be numerically modeled. This paper is organized as follows. Section II describes the model used. Section III presents the possible stationary states of the nanobubbles. The stability of these stationary states against thermal fluctuations is analyzed in Section IV. The main results are summarized in the Conclusion (Section V).
II Model
[…]
Consider a rectangular graphene sheet in which the chains of valent bonds along the $x$-axis have a ”zigzag” structure with $N_{x}$ nodes, and along the $y$-axis—an ”armchair” structure with $N_{y}$ nodes (the sheet lies parallel to the $z=0$ plane). Let us take a sheet with free edges—see Fig. 1. The sheet consists of $N_{c}=N_{x}N_{y}/2-2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its length is $L_{x}=(N_{x}-1)r_{c}\sqrt{3}/2$ and its width is $L_{y}=(3N_{y}/4-1)r_{c}$ ($r_{c}=1.418$ Åis the C–C bond length). In all calculations, we will use a graphene sheet with $N_{x}=416$, $N_{y}=480$ (sheet dimensions: $50.963\times 50.906$ nm2, number of carbon atoms $N_{c}=99838$). We will assume that hydrogen atoms are attached to the edge carbon atoms, which participate in the formation of only two C–C bonds, forming united CH atoms with them of mass $M_{e}=13m_{p}$ (the mass of the internal sheet atoms is $M_{c}=12m_{p}$, where $m_{p}=1.66\times 10^{-27}$ kg is the proton mass).
Figure 1: A rectangular graphene sheet with $N_{x}=10$, $N_{y}=12$. Black disks represent carbon atoms C, black solid lines represent valence bonds C–C. Small light disks indicate the positions of the centers of mass of the valence bond hexagons. Dashed lines illustrate the triangulation of the graphene sheet using these centers – the inner surface of the sheet is divided into triangles, each containing one carbon atom (the triangle corresponding to the $n$-th atom is highlighted; $\mathbf{u}_{n}$ is the coordinate vector of the atom).
The Hamiltonian of the nanosheet is
${\cal H}_{1}=\sum_{n=1}^{N_{c}}\left[\frac{1}{2}M_{c}(\dot{\bf u}_{n},\dot{\bf u}_{n})+Q_{n}+Z({\bf u}_{n})\right],$ (1)
where the vector $\mathbf{u}_{n}=(x_{n}(t),y_{n}(t),z_{n}(t))$ defines the coordinates of the $n$-th carbon atom at time $t$. The first term of the sum in equation (1) represents the kinetic energy, while the second term $Q_{n}$ describes the valence interaction of the atom with index $n$ with other atoms in the graphene sheet. This includes deformations of valence bonds, valence angles, and torsion angles—a detailed description of the force field used is provided in Savin2010 .
The valence bond between two neighboring carbon atoms $n$ and $k$ can be described by the Morse potential
$U_{1}({\bf u}_{n},{\bf u}_{k})=\epsilon_{c}\left\{\exp[-\alpha(r_{nk}-r_{c})]-1\right\}^{2},$ (2)
where $r_{nk}=|\mathbf{u}_{n}-\mathbf{u}_{k}|$ is the distance between the atoms, and the bond energy is $\epsilon_{c}=4.9632$ eV.
Valence angle deformation energy between three adjacent carbon atoms $n$, $k$, and $l$ can be described by the potential
$U_{2}({\bf u}_{n},{\bf u}_{k},{\bf u}_{l})=\epsilon_{\varphi}(\cos\varphi-\cos\varphi_{0})^{2},$ (3)
where $\cos\varphi=({\bf u}_{n}-{\bf u}_{k},{\bf u}_{l}-{\bf u}_{k})/r_{nk}r_{kl}$, and $\varphi_{0}=2\pi/3$ is the equilibrium valent angle. Parameters $\alpha=1.7889$ nm-1 and $\epsilon_{\varphi}=1.3143$ eV can be found from the small amplitude oscillations spectrum of the graphene sheet Savin2008 .
Valence bonds between four adjacent carbon atoms $n$, $m$, $k$ , and $l$ constitute torsion angles, the potential energy of which can be defined as
$U_{3}({\bf u}_{n},{\bf u}_{m},{\bf u}_{k},{\bf u}_{l})=\epsilon_{\phi}(1-\cos\phi),$ (4)
where $\phi$ is the corresponding torsion angle ($\phi=0$ is the equilibrium value of the angle) and $\epsilon_{\phi}=0.499$ eV.
The last term in the sum (1) describes the interaction of the sheet atoms with the flat substrate on which it rests. Let the flat substrate occupy the half-space $z\leq 0$. In this case, the interaction energy of a carbon atom with the substrate can be described using a (3,9) Lennard-Jones potential Aitken2010 ; Zhang2013 ; Zhang2014 :
$Z({\bf u})=Z(z)=\varepsilon_{0}\left[(h_{0}/z)^{9}-3(h_{0}/z)^{3}\right]/2,$ (5)
where $\varepsilon_{0}$ is the interaction energy (adhesion energy) and $h_{0}$ is the equilibrium distance to the surface of the half-space. For a silicon dioxide (SiO2) substrate, the energy is $\varepsilon_{0}=0.074$ eV and the distance is $h_{0}=5$ Å  Koenig2011 .
Table 1: Values of the parameters of the interaction potential with a flat substrate (5) for atoms C, Ar, Kr, Xe.
C Ar Kr Xe
$\varepsilon_{0}$ (eV)  0.074  0.1406  0.17242  0.19462
$h_{0}$ (Å) 5.0 5.0 5.116 5.397
The Hamiltonian of the inert gas molecular subsystem has the form
${\cal H}_{2}=\sum_{n=N_{c}+1}^{N}\left[\frac{1}{2}M_{g}(\dot{\bf u}_{n},\dot{\bf u}_{n})+Z({\bf u}_{n})\right]+E_{2},$ (6)
where $N=N_{c}+N_{g}$ is the total number of atoms in the two-component molecular system, the vectors $\{{\bf u}_{n}\}_{n=N_{c}+1}^{N}$ specify the coordinates of the gas atoms, and $M_{g}$ is the atomic mass ($M_{g}=39.96$, 83.80, 131.29 $m_{p}$ for Ar, Kr, and Xe atoms, respectively). The first term in the sum (6) represents the kinetic energy of the gas subsystem, the second term—the interaction energy of the gas atoms with the flat substrate (the values of the interaction potential parameters (5) for the gas atoms are given in Table 1). The last term,
$E_{2}=\sum_{n=N_{c}+1}^{N-1}\sum_{k=n+1}^{N}V(r_{nk}),$ (7)
defines the interaction energy between the gas atoms, where $r_{nk}=|{\bf u}_{n}-{\bf u}_{k}|$ is the distance between atoms $n$ and $k$. The interaction between atoms is described by the Lennard-Jones (6,12) potential:
$V(r)=4\epsilon[(\sigma/r)^{12}-(\sigma/r)^{6}].$ (8)
The values of the interaction potential parameters for various pairs of atoms are given in Table 2.
Table 2: Values of the interaction potential parameters (8) for pairs of atoms CAr, CKr, CXe, ArAr, KrKr, XeXe Iakovlev2017 ; Rappe1992 ; Vogt2014 .
CAr CKr CXe ArAr KrKr XeXe
$\varepsilon_{1}$ (eV)  0.00525  0.00643  0.00724  0.010  0.015  0.019
$\sigma$ (Å) 3.3965 3.4965 3.7465 3.40 3.60 4.10
The interaction energy between the gas atoms and the graphene sheet is given by the sum
$E_{3}=\sum_{n=1}^{N_{c}}\sum_{k=N_{c}+1}^{N}V(r_{nk}).$ (9)
The total Hamiltonian of the two-component molecular system is
$\mathcal{H}=\mathcal{H}_{1}+\mathcal{H}_{2}+E_{3}.$ (10)
The potential energy of the molecular system is
$E=\sum_{n=1}^{N_{c}}Q_{n}+\sum_{n=1}^{N}Z(\mathbf{u}_{n})+E_{2}+E_{3}.$ (11)
Figure 2: Stationary $l$-layer states of encapsulated $N_{g}=2909$ argon atoms with $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). For clarity of presentation, only one half of the graphene sheet is shown. The graphene sheet is depicted in gray. For argon atoms, their vertical displacement is indicated by color (blue corresponds to the minimum displacement with coordinate $z=h_{0}$, red corresponds to the maximum displacement with $z=h_{0}+10$ Å).
Figure 3: Distribution of vertical displacements $p(z)$ for the stationary states of encapsulated $N_{g}=3936$ krypton atoms with number of layers $l=1$, …, 5.
More detailed discussion and motivation of our choice of the interaction potentials (2), (3), (4) can be found in Savin2010 . Note that in molecular dynamics of carbon nanotubes and nanoribbons, more complex reactive potentials are often used, such as Tersoff Lindsay2010 , REBO Stuart2000 ; Brener2002 and ReaxFF Srinivasan2015 . These potentials are able to describe the formation and breaking of bonds between carbon atoms, but they are worse in describing the frequency spectrum of graphene and are more complicated for numerical modeling. The use of these potentials is justified only when modeling processes in graphene-based systems that occur with changes in the topology of the valence bond lattice. Nevertheless, these potentials reliably enough describe elastic properties of graphene Lebedeva2019 , so the decision on use them will lead to the same results.
III Stationary States of Graphene Nanobubbles
To find a stationary state of an encapsulated molecular system (nanobubble), it is necessary to numerically solve the potential energy minimization problem
$E\rightarrow\min:\{{\bf u}_{n}\}_{n=1}^{N}.$ (12)
The problem (12) was solved numerically using the conjugate gradient method Fletcher1964 ; Shanno1976 .
As initial configurations, structures in the form of an $l$-layer stepped pyramid of gas atoms ($l=1,2,\ldots$) covered from above by a graphene sheet were used. By varying the initial configuration, all possible stable stationary states of the encapsulated molecular system can be obtained.
Let $\{{\bf u}_{n}^{0}=(x_{n}^{0},y_{n}^{0},z_{n}^{0})\}_{n=1}^{N}$ be a solution to problem (12). The corresponding stationary state is characterized by its energy $E(\{{\bf u}_{n}^{0}\}_{n=1}^{N})$, the maximum value of the vertical displacements of the graphene sheet over the encapsulated group of atoms (the height of the graphene bulge)
$H=\max_{n=1,\ldots,N_{c}}(z_{n}^{0}-h_{0}),$
the radius $R$ of the bulge boundary, as well as the density $d$ and pressure $P$ inside the nanobubble.
We will consider that a sheet atom participates in forming the upper surface of the nanobubble if its vertical displacement $\Delta z_{n}=z_{n}^{0}-h_{0}>h_{1}$, where the displacement threshold is $h_{1}=1.5$ Å. The center of the nanobubble then has coordinates
$x_{c}=\frac{1}{N_{i}}\sum_{n=1\atop\Delta z_{n}>h_{1}}^{N_{c}}x_{n}^{0},\quad y_{c}=\frac{1}{N_{i}}\sum_{n=1\atop\Delta z_{n}>h_{1}}^{N_{c}}y_{n}^{0},$
where the number of graphene sheet atoms forming the upper surface of the bubble is
$N_{i}=\sum_{n=1\atop\Delta z_{n}>h_{1}}^{N_{c}}1.$
It is convenient to define the base radius of the bubble as
$R=\frac{1}{N_{r}}\sum_{n=1\atop\Delta z_{n}\in[h_{1},h_{2}]}^{N_{c}}r_{n},\quad N_{r}=\sum_{n=1\atop\Delta z_{n}\in[h_{1},h_{2}]}^{N_{c}}1.$
Here, the second vertical displacement threshold is $h_{2}=2.5$ Å, and the distance to the center is $r_{n}=[(x_{n}^{0}-x_{c})^{2}+(y_{n}^{0}-y_{c})^{2}]^{1/2}$.
The volume of the encapsulation cavity beneath the graphene sheet (the nanobubble volume) is
$V=\sum_{n=1}^{N_{c}}\Delta z_{n}S_{n}^{0},$
where $S_{n}^{0}$ is the area of the projection onto the plane $z=0$ of the triangle corresponding to the $n$-th sheet atom (see the graphene sheet triangulation scheme in Fig. 1). The density of gas atoms inside the nanobubble is $d=M_{g}N_{g}/V$.
Figure 4: Stationary $l$-layer states of a nanobubble with $N_{g}=3936$ krypton atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of deformation energy across the graphene sheet (blue corresponds to zero deformation, red corresponds to deformation with an energy of 0.008 eV per atom).
Figure 5: Stationary $l$-layer states of a nanobubble with $N_{g}=2909$ xenon atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e). A top view is shown. Color indicates the distribution of pressure across the graphene sheet (blue corresponds to zero pressure, red corresponds to a pressure of $p_{n}=5$ GPa). The average pressure values are $P=1.09$, 0.90, 0.86, 0.84, and 0.85 GPa.
In the stationary state, a force acts on the $n$-th atom of the graphene sheet from the side of the encapsulated atoms:
${\bf F}_{n}=\frac{4\epsilon_{0}}{\sigma^{2}}\sum_{k=N_{c}+1}^{N}\left[2(\sigma/r_{nk})^{14}-(\sigma/r_{nk})^{8}\right]({\bf u}_{n}^{0}-{\bf u}_{k}^{0}),$
where the distance between atoms is $r_{nk}=|{\bf u}_{n}^{0}-{\bf u}_{k}^{0}|$. Consequently, the local pressure on this atom is $p_{n}=({\bf F}_{n},{\bf e}_{n})/S_{n}$, where ${\bf e}_{n}$ is the unit vector orthogonal to the plane of the triangle containing the sheet atom, and $S_{n}$ is the area of this triangle (see Fig. 1). The value of the average pressure on the external surface of the nanobubble is
$P=\frac{1}{N_{i}}\sum_{n=1\atop\Delta z_{n}>h_{1}}^{N_{c}}p_{n}.$
This value also determines the average value of the internal pressure.
Table 3: Dependence on the number of encapsulated argon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble. Values for the ground states are highlighted in italics.
$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
188 1 0 1.428 3.23 25.33 1.70 0.128
2 0.0188 1.307 5.90 22.98 1.64 0.128
260 1 0 1.257 3.23 29.84 1.76 0.108
2 0.0137 1.302 6.12 26.33 1.68 0.232
398 1 0 1.171 3.23 36.71 1.80 0.088
2 0.0063 1.149 6.23 31.38 1.74 0.199
614 1 0 1.080 3.23 45.71 1.85 0.071
2 0.0015 1.030 6.22 37.64 1.79 0.165
1 0 1.012 3.23 57.33 1.87 0.056
962 2 -0.0018 0.950 6.22 46.57 1.82 0.165
3 0.0010 0.953 9.11 44.30 1.79 0.206
1 0 0.974 3.23 67.34 1.90 0.048
1335 2 -0.0046 0.866 6.22 53.08 1.85 0.117
3 -0.0046 0.892 9.24 50.68 1.80 0.183
1 0 0.942 3.23 78.83 1.91 0.041
1828 2 -0.0063 0.801 6.22 61.26 1.87 0.102
3 -0.0048 0.793 9.23 56.41 1.83 0.164
4 -0.0003 0.796 12.02 54.32 1.81 0.221
1 0 0.906 3.24 99.46 1.93 0.033
2 -0.0083 0.741 6.23 75.84 1.89 0.082
2909 3 -0.0063 0.767 9.25 73.10 1.85 0.127
4 -0.0065 0.708 12.20 66.67 1.83 0.183
5 -0.0032 0.705 14.55 64.18 1.82 0.227
1 0 0.885 3.24 115.70 1.94 0.028
2 -0.0084 0.723 6.24 88.75 1.90 0.070
3936 3 -0.0084 0.723 9.24 82.77 1.86 0.112
4 -0.0097 0.696 12.24 76.09 1.84 0.161
5 -0.0079 0.678 15.02 73.79 1.83 0.204
Table 4: Dependence on the number of encapsulated krypton atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
188 1 0 1.442 3.35 27.46 3.02 0.122
2 0.0234 1.509 5.99 25.78 2.86 0.232
260 1 0 1.357 3.35 32.23 3.11 0.084
2 0.0179 1.389 6.45 28.48 2.96 0.226
398 1 0 1.284 3.35 39.91 3.17 0.084
2 0.0086 1.240 6.53 33.90 3.05 0.193
1 0 1.194 3.36 49.24 3.26 0.068
614 2 0.0041 1.133 6.52 40.75 3.13 0.160
3 0.0145 1.151 9.07 39.78 3.06 0.228
1 0 1.131 3.35 61.63 3.30 0.054
962 2 -0.0006 1.063 6.53 49.85 3.19 0.131
3 0.0068 1.023 9.61 47.47 3.12 0.202
1 0 1.098 3.35 72.43 3.35 0.046
1335 2 -0.0033 0.982 6.53 57.02 3.24 0.115
3 0.0015 0.964 9.72 53.34 3.16 0.182
4 0.0930 0.959 12.49 51.15 3.14 0.206
1 0 1.069 3.35 84.72 3.37 0.040
1828 2 -0.0055 0.944 6.54 66.20 3.26 0.099
3 -0.0029 0.892 9.70 60.17 3.20 0.161
4 0.0005 0.941 12.41 60.66 3.17 0.205
1 0 1.038 3.35 106.65 3.41 0.031
2 -0.0080 0.872 6.54 81.66 3.31 0.080
2909 3 -0.0070 0.852 9.71 74.79 3.23 0.130
4 -0.0047 0.813 12.94 71.05 3.19 0.182
5 -0.0015 0.877 15.41 72.16 3.18 0.214
1 0 1.018 3.36 123.73 3.44 0.027
2 -0.0096 0.827 6.55 93.70 3.34 0.070
3936 3 -0.0094 0.806 9.71 85.27 3.25 0.114
4 -0.0082 0.761 12.91 79.40 3.21 0.163
5 -0.0061 0.796 15.98 78.45 3.20 0.204
The solution of problem (12) showed that each stationary state of a nanobubble always has a strictly defined number of atomic layers, $l$, in the cluster of internal atoms—see Figs. 2 and 3. For a number of encapsulated atoms $N_{g}<4000$, stationary states with $l=1,\ldots,6$ layers are possible. Each internal atom lies within one of the layers (see Fig. 3). The layers have a flat, circular shape (ranging from the shape of a regular hexagon to a circle). The layers are concentrically stacked on top of each other and form a symmetric $l$-stepped pyramid with a flat top (the plane of the layers is always parallel to the substrate plane). Atoms within each layer are in the same plane, with only edge atoms displaced downwards, smoothing the stepped shape of the pyramid. Atoms Ar, Kr, and Xe differ in size, hence the interlayer distances for them are $\Delta h=3.0$, 3.2, and 3.6 Å, respectively. For argon atoms, two- and three-layer structures at $N_{g}=2463$ were first obtained in Iakovlev2017 .
Covering the $l$-stepped pyramid from above with a graphene sheet is achieved through its local stretching. The valence bonds of the sheet stretch over the group of encapsulated atoms, while outside this zone the sheet remains in an undeformed state, lying flush against the substrate—see Fig. 4. The most significant stretching of C–C bonds occurs over the layer boundaries, where the deformation energy can reach values of 0.01 eV per carbon atom. As seen in Fig. 4, the zones of sheet stretching strictly correspond to the stepped structure of the pyramid of encapsulated atoms. For example, for $N_{g}=3936$ and $l$-layer states with $l=1,2,3$, stretching of the sheet occurs only over the edges of the layers. For states with a larger number of layers ($l=4,5$), stretching occurs more uniformly across the entire surface of the nanobubble, but the maximum values are always reached at its base (over the edge of the first layer).
Local stretching of the valence bonds of the graphene sheet induces pressure $p_{n}$ on the encapsulated atoms. As shown in Fig. 5, the pressure arising from encapsulation is localized only on the surface of the nanobubble. Over the layer boundaries, the pressure can reach maximum values of up to 10 GPa, while the average pressure value is $P\sim 1$ GPa. The highest pressure is experienced by single-layer packings of internal atoms. The average pressure value decreases with an increasing number of layers.
Table 5: Dependence on the number of encapsulated xenon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
188 1 0 1.476 3.57 31.54 3.52 0.113
2 0.0278 1.492 6.87 28.87 3.28 0.231
260 1 0 1.361 3.58 36.85 3.60 0.097
2 0.0216 1.367 7.15 32.62 3.34 0.219
398 1 0 1.296 3.58 45.58 3.66 0.079
2 0.0061 1.096 7.21 38.37 3.45 0.188
1 0 1.227 3.59 56.15 3.75 0.064
614 2 -0.0009 1.133 7.23 45.86 3.50 0.158
3 0.0108 1.165 10.36 44.91 3.41 0.213
1 0 1.180 3.58 70.11 3.80 0.051
962 2 -0.0068 1.083 7.23 56.27 3.55 0.128
3 -0.0013 1.015 10.76 53.26 3.45 0.202
1 0 1.146 3.58 82.44 3.84 0.043
1335 2 -0.0117 1.006 7.24 65.08 3.57 0.111
3 -0.0104 0.965 10.88 60.08 3.49 0.181
4 00.0012 0.980 14.04 58.02 3.45 0.242
1 0 1.125 3.59 96.27 3.86 0.037
1828 2 -0.0160 0.953 7.25 74.47 3.62 0.097
3 -0.0306 0.880 10.93 67.94 3.51 0.161
4 -0.0124 0.925 14.36 65.59 3.48 0.219
1 0 1.090 3.60 120.99 3.90 0.030
2 -0.0198 0.897 7.27 92.58 3.65 0.079
2909 3 -0.0238 0.860 10.89 83.19 3.55 0.131
4 -0.0228 0.841 14.55 79.27 3.50 0.184
5 -0.0171 0.846 17.81 76.76 3.47 0.232
1 0 1.077 3.60 140.68 3.91 0.026
2 -0.0218 0.872 7.26 106.36 3.67 0.068
3936 3 -0.0272 0.828 10.88 95.06 3.56 0.114
4 -0.0286 0.791 14.53 87.11 3.52 0.167
5 -0.0255 0.794 17.79 88.19 3.47 0.202
6 -0.0233 0.804 20.81 85.72 3.48 0.243
The maximum possible number of layers, $l_{m}$, depends on the number of internal atoms, $N_{g}$ ($l_{m}$ increases with increasing $N_{g}$). For argon atoms, only single-layer packings ($l_{m}=1$) are possible for $N_{g}\leq 150$, while for $150<N_{g}<900$, single- and double-layer packings ($l_{m}=2$) are possible. For $900<N_{g}<1800$, the maximum number of layers is $l_{m}=3$; for $1800<N_{g}<2800$, it is $l_{m}=4$; and for $2800<N_{g}<4000$, it is $l_{m}=5$. A single-layer packing exists for any number of atoms. Therefore, the relative energy of a stationary state can be measured from the energy of the single-layer packing: $\Delta E=(E_{l}-E_{1})/N_{g}$, where $E_{l}$ is the energy of the $l$-layer stationary state.
The dependence of the relative specific energy $\Delta E$, internal pressure $P$, height $H$, radius $R$, and density $d$ of the nanobubble’s stationary state on the number of argon atoms $N_{g}$ is given in Table 3. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$. For $900<N_{g}<3000$, two-layer states become more favorable, and for $N_{g}>3000$, three- and four-layer states are more favorable. The pressure inside the encapsulation monotonically decreases with an increasing number of atoms. With an increasing number of layers, a slight decrease in pressure and density is observed. The height of the nanobubble, $H$, always takes a discrete set of values proportional to the number of layers $l$: $H\approx h_{0}+(l-1)\Delta h$.
For krypton atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of internal atoms $N_{g}$ is presented in Table 4. Here, for $N_{g}\leq 140$, the maximum number of layers is $l_{m}=1$; for $140<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$, while two-layer states are favorable for $900<N_{g}<4000$. Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms. Pressure and density reach their maximum values for single-layer states.
For xenon atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on $N_{g}$ is presented in Table 5. Here, for $N_{g}\leq 116$, the maximum number of layers is $l_{m}=1$; for $116<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; for $2800<N_{g}<3800$, $l_{m}=5$; and for $3800<N_{g}<4000$, $l_{m}=6$. As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<600$, two-layer states are favorable for $600<N_{g}<1400$, and three-layer states are favorable thereafter. Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms, and pressure and density reach their maximum values for single-layer packings.
Figure 6: Dependence on temperature $T$ of (a) height $H$, (b) density $d$, and (c) internal pressure $P$ for a graphene nanobubble with $N_{g}=2909$ xenon atoms. Curves 1, 6, 11 correspond to the state with $l=1$; curves 2, 7, 12 – to $l=2$, …, and curves 5, 10, 15 – to $l=5$ layers. The dashed vertical lines correspond to the transition temperatures between states $l\rightarrow 4$: $T_{1}=960$, $T_{2}=390$, $T_{3}=330$, and $T_{5}=210$ K. The dotted vertical lines correspond to the melting temperature of the single-layer (420 K) and four-layer state (270 K).
Figure 7: Dependence on temperature $T$ of the vertical displacement distribution $p(z)$ for $l$-layer states of $N_{g}=2909$ internal xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
IV Stability of Stationary States to Thermal Fluctuations
Note that all stationary states of nanobubbles obtained by solving the energy minimization problem are always stable. To verify their stability against thermal fluctuations, we performed simulations of nanobubble dynamics at temperatures $T\leq 1080$ K.
Figure 8: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
Figure 9: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
Dynamics of a thermalized graphene sheet with encapsulated atoms is described by the system of Langevin equations:
$M_{n}\ddot{\bf u}_{n}=-\frac{\partial H}{\partial{\bf u}_{n}}-\Gamma M_{n}\dot{\bf u}_{n}+\Xi_{n},~~n=1,...,N,$ (13)
where the mass $M_{n}=M_{c}$ for $n=1,...,N_{c}$ and $M_{n}=M_{g}$ for $n=N_{c}+1,...,N$ (the total number of atoms is $N=N_{c}+N_{g}$). Here, $\Gamma=1/t_{r}$ is the friction coefficient characterizing the intensity of energy exchange with the thermostat (relaxation time $t_{r}=10$ ps), and $\Xi_{n}=\{\xi_{n,i}\}_{i=1}^{3}$ is a three-dimensional vector of normally distributed random forces normalized by the following conditions:
$\langle\xi_{n,i}(t)\xi_{k,j}(s)\rangle=2\Gamma M_{n}k_{B}T\delta_{nk}\delta_{ij}\delta(t-s).$
Here, $T$ is the thermostat temperature and $k_{B}$ is the Boltzmann constant.
The relaxation time $t_{r}$ determines the intensity of energy exchange with the thermostat. In the simulated molecular system, the role of the thermostat is played by a flat substrate, to which the atoms are weakly coupled via non-valence interactions. For such interactions, the relaxation time is typically $t_{r}\sim 100$ ps. For the convenience of numerical simulation, we use a smaller value, $t_{r}=10$ ps. This allows us to significantly reduce the numerical integration time required to reach an equilibrium state and obtain reliable averages.
We numerically integrate the system of equations of motion (13) with initial conditions corresponding to a stationary nanobubble state:
$\{{\bf u}_{n}(0)={\bf u}_{n}^{0},~\dot{\bf u}_{n}(0)=0\}_{n=1}^{N}.$ (14)
The equations of motion Eq. (13) with initial conditions (14) are solved numerically using the velocity Verlet method Verlet1967 . A time step of 1 fs is used in the simulations, as further reduction of the time step has no appreciable effect on the results.
Numerical simulations of the dynamics showed that among the $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases. All other states transition into this ground state at a specific temperature. The dependence of the nanobubble’s height $H$, density $d$, and internal pressure $P$ on temperature is shown in Fig. 6.
[…]
In Tables 3, 4, 5, the rows corresponding to the ground states are given in italics. Note that the ground states are not always the most energetically favorable. Their key feature is the smooth (non-abrupt) change in their shape with increasing temperature.
Conformational changes of the encapsulated atom cluster are reflected in the change in the distribution of atoms by their vertical $\{z_{n}\}_{n=N_{c}}^{N}$ (Fig. 7) and radial (in-plane) displacements from the center of mass:
$\{r_{n}=[(x_{n}-x_{c})^{2}+(y_{n}-y_{c})^{2}]^{1/2}\}_{n=N_{c}+1}^{N}$
(Figs. 8, 9).
As can be seen from Fig. 7 (a, b, c), the shape of the vertical displacement distribution changes sharply at the temperature of $l$-layer state stability loss. For $l\neq 4$, the clearly pronounced layered structure of the cluster disappears here. For the ground state ($l=4$), only a gradual, smooth broadening of the layered structure occurs with increasing temperature, with only the first layer of atoms adjacent to the substrate remaining clearly pronounced—see Fig. 7 (d).
Radial displacements also show changes in the distribution shape at the transition temperature $T_{l}$. The transition to a state with a larger number of layers leads to a sharp decrease in the diameter of the inner atom cluster—see Figs. 8 (a, b, c) and 9 (a, b, c). For the ground state with $l=4$ layers, only a continuous, gradual increase in the cluster diameter (its thermal expansion) occurs—see Figs. 8 (d) and 9 (d).
The shape of the radial distribution also allows one to judge the phase state of the cluster. At low temperature, the radial distribution has a discrete structure characteristic of a crystalline (solid) state. Upon reaching the melting temperature $T_{l,m}$, the discrete distribution shape is replaced by a continuous one, characteristic of a liquid state. For $N_{g}=2909$ xenon atoms, the melting of the single-layer cluster occurs at $T_{1,m}=420\div 450$ K, and of the four-layer cluster at $T_{4,m}=270\div 300$ K—see Fig. 8. For argon atoms, the melting temperatures are $T_{1,m}=240\div 270$ K, $T_{4,m}=150\div 180$ K (see Fig. 9), and for krypton $T_{1,m}=300\div 330$ K, $T_{4,m}=210\div 240$ K. Such high melting temperatures are explained by the high value of internal pressure—see Fig. 6 (c).
Figure 10: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble: with $N_{g}=962$ and 2909 argon atoms (curves 1 and 2); with $N_{g}=3936$ krypton atoms (curve 3); with $N_{g}=2909$ xenon atoms (curve 4).
The existence of several stable states with different numbers of atom layers $l$ at low temperatures indicates that the nanobubbles here cannot have a universal (scalable) shape, and the height-to-radius ratio is not constant. Indeed, for a single-layer state ($l=1$), increasing the number of internal atoms $N_{g}$ does not change the nanobubble’s height ($H\approx 3.3$ Å, while its radius $R$ grows as $\sqrt{N_{g}}$; therefore, the ratio $H/R\rightarrow 0$ as $N_{g}\rightarrow\infty$. Based on the data in Tables 3, 4, and 5, the ratio $H/R$ can vary from 0 to 0.24 here. The violation of the scaling rule for graphene bubbles with a radius of $\sim 1$ nm was noted in the experimental study Villarreal2021 , where they were created via the implantation of noble gas ions He, Ne, and Ar. As can be seen from the table data, the relation $H/R\approx 0.2$ holds only for the ground states of the nanobubble.
The dependence of the ratio $H/R$ on temperature $T$ for the ground states of the nanobubble is shown in Fig. 10. As can be seen from the figure, the value of $H/R$ increases monotonically with temperature, while the spread of values decreases: $H/R$ varies within the intervals [0.163, 0.206] at $T=0$ K; [0.192, 0.208] at $T=300$ K; [0.211, 0.217] at $T=600$ K. These estimates are in good agreement with the universal value $H/R=0.204$ obtained in Jain2017 for a graphene nanobubble with an ideal gas. Note that the study Korneva2023 reported an estimate of $H/R\approx 0.13$ for graphene nanobubbles with argon atoms at $T=2$ K, indicating that the simulation used not a ground stationary state of the nanobubble, but rather a state with a smaller number of layers.
V Conclusion
Using the example of inert gas atoms (Ar, Kr, Xe), it has been shown that graphene nanobubbles located on flat substrates are multistable systems. A nanobubble can have many stable stationary states, each characterized by the number of layers $l$ in the internal atom cluster. Inside the nanobubble, each atom lies within one of the layers. The layers have a circular, flat shape, are concentrically stacked on top of each other, and form an $l$-stepped pyramid with a flat top. Atoms within each layer lie in a single plane, with only the edge atoms displaced downwards, smoothing the stepped shape of the pyramid. The covering of the pyramid from above by a graphene sheet occurs due to its local stretching. The sheet’s valence bonds are stretched only above the group of encapsulated atoms; outside this coverage zone, the sheet remains undeformed and tightly adheres to the substrate. The strongest stretching (up to one percent) occurs above the edges of the layers.
The maximum possible number of layers $l_{m}$ increases monotonically with the number of internal atoms $N_{g}$. For argon atoms, only single-layer stationary states ($l_{m}=1$) are possible for $N_{g}\leq 150$; for $150<N_{g}<900$—single-layer and double-layer states ($l_{m}=2$); for $900<N_{g}<1800$, $l_{m}=3$; for $1800<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$. The graphene sheet, due to van der Waals interaction with the substrate, presses the cluster of encapsulated atoms to substrate, creating pressures $P\sim 1$ GPa. The highest pressure is experienced by single-layer packings of internal atoms. The magnitude of the pressure decreases as the number of layers increases.
Thermal fluctuation modeling shows that among all $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases. The other states transition into this state at a specific temperature $T_{l}$. For example, for $N_{g}=2909$ encapsulated atoms, the ground state is the four-layer packing ($l=4$). For Xe atoms, the $l$-layer states persist up to temperatures $T_{l}=960$, 390, 330, 210 K; for Ar atoms, $T_{l}=600$, 270, 180, 30 K; and for Kr atoms—750, 360, 240, 150 K, where $l=1$, 2, 3, 5.
The existence of several stable states with different numbers of layers in the internal atom cluster at low temperatures indicates that universal scaling does not hold for these nanobubbles. The height-to-radius ratio $H/R$ of the nanobubble is not constant and can vary from 0 to 0.24. The universal relation $H/R\approx 0.2$ holds only for the ground states.
Acknowledgements
The research was funded by the Russian Science Foundation (RSF) (project No. 25-73-20038).

### v3

@@
 Abstract
-Using the example of Ar, Kr, and Xe atoms, it is shown that graphene nanobubbles on flat substrates are multistable systems.
-A nanobubble can have many stable stationary states, each characterized by the number of layers, $l$, within the cluster of internal atoms.
-The layers are circular in shape, concentrically stacked on top of each other, forming an $l$-stepped pyramid with a flat top.
-The covering of this pyramid with a graphene sheet is achieved through its local stretching.
-The valence bonds of the sheet stretch only over the group of internal atoms; outside the coverage zone, the sheet remains undeformed and lies flush against the substrate.
-The maximum possible number of layers, $l_{m}$, increases monotonically with the number of atoms $N$ ($l_{m}=6$ for $N=4000$).
-The graphene sheet, interacting with the substrate, compresses the internal atom cluster against it, generating an internal pressure of $P\sim 1$ GPa.
-Numerical simulations of thermal vibrations reveal that among all $l$-layer configurations of a nanobubble, there is always one ”ground” state.
-Upon heating, this ground state smoothly transitions into a layerless liquid state.
-All other stationary states transform into this ground state once a certain temperature is reached (for $N=4000$, the ground state corresponds to state with $l=4$).
-The coexistence of several stable states with different numbers of layers at low temperatures leads to the absence of a universal shape for the nanobubbles.
-In this scenario, the height-to-radius ratio, $H/R$, can vary from 0 to 0.24, depending on the number of layers.
-Keywords: Van der Walls bubble, graphene nanobubble, molecular modelling, phase transition
+Using He, Ne, Ar, Kr, and Xe atoms as a model system, it is demonstrated that graphene nanobubbles on flat substrates are multistable systems.
+A nanobubble can adopt multiple stable stationary states, each characterized by the number of layers $l$ within the cluster of encapsulated atoms.
+The layers are circular, concentrically stacked, and form an $l$-stepped pyramid with a flat top.
+Encapsulation of this pyramid by the graphene sheet is achieved through local stretching of the membrane: the valence bonds elongate only directly above the confined atoms.
+Outside this coverage zone, the sheet remains undeformed and lies flush against the substrate.
+The maximum number of possible layers, $l_{m}$, increases monotonically with the number of encapsulated atoms $N$, reaching $l_{m}=6$ for $N=4000$.
+The graphene membrane, through van der Waals interaction with the substrate, compresses the internal atomic cluster, generating pressures on the order of $P\sim 1$ GPa.
+Numerical simulations of thermal vibrations reveal that among all $l$-layer configurations, one ground state always exist.
+Upon heating, this state smoothly transitions into a layerless liquid configuration.
+All other stationary states transform into this ground state once a characteristic temperature $T_{l}$ is reached.
+For $N=4000$, the ground state corresponds to the four-layer packing ($l=4$).
+The coexistence of multiple stable states with distinct layer numbers at low temperatures leads to the absence of a universal shape for the nanobubbles.
+In this scenario, the height-to-radius ratio, $H/R$ is not constant and can vary from 0 to 0.28, depending on the number of layers.
+Keywords: Van der Walls encapsulation, graphene nanobubble, molecular modelling, phase transition, aspect ratio
 I Introduction
@@
 In graphene nanobubbles, high pressure is induced by the forces of interaction between the sheet and the substrate.
+Graphene nanobubbles have the potential to become indispensable tools in nanotechnology.
+Analysis of their shape allows for the determination of the adhesion energy between a graphene sheet and a substrate Wang2016 .
+Nanobubbles in multilayer graphene can be utilized in the development of cathode materials for lithium-ion batteries He2023 .
+Furthermore, they can be used to create two-dimensional van der Waals solids, which may find applications in quantum information technologies Langle2024 .
 Encapsulation structures (nanobubbles) can form spontaneously as surface defects during the mechanical transfer of two-dimensional sheets in the creation of van der Waals heterostructures Khestanova2016 ; Sanchez2018 .
 Nanobubbles can also be intentionally fabricated using ion bombardment Blundo2020 ; Zamborlini2015 ; Villarreal2021 or selective adsorption methods Zahra2020 .
-In addition to experimental methods, analytical models and numerical simulations are used to study nanobubbles.
-Analytical models employ elastic plate theory, taking into account the contribution of adhesion energy between the layers of van der Waals heterostructures Yue2012 ; Wang2013a ; Dai2018 .
-More complex models also incorporate the equation of state of the substance inside the nanobubble Zhilyaev2019 ; Iakovlev2019 ; Aslymov2020 ; Qu2023 and nonlinear effects Lyublinskaya2020 .
-Using such models, it has been shown that nanobubbles with a circular shape exhibit universal scaling behavior: the ratio of the nanobubble’s height to its radius, $H/R$, remains constant.
-Numerical simulations allow not only for the verification of analytical results but also for gaining deeper insights into nanobubble properties.
-It has been demonstrated that spatial confinement and internal pressure within a nanobubble promote the formation of internal crystalline structures, which can be planar in shape Iakovlev2017 , and that the nanobubble’s shape can critically depend on the properties of the encapsulated substance Ghorbanfekr2017 .
-Numerical simulations have shown Jain2017 that a nanobubble containing an ideal gas always has a circular shape with a constant ratio $H/R=0.204$.
-Molecular dynamics simulations have modeled transitions of argon inside a nanobubble into a liquid-crystal state Korneva2023 .
-It has been demonstrated that the presence of nanobubbles can lead to a reduction in the interfacial thermal conductivity of layered van der Waals structures Qu2023a , as well as cause their significant thermal contraction Qu2023b .
-In the paper, all possible stationary states of graphene nanobubbles will be numerically modeled.
-This paper is organized as follows.
-Section II describes the model used.
+In addition to experimental techniques, analytical models and numerical simulations have become essential tools for investigating nanobubbles.
+Analytical approaches are often based on elastic plate theory, which incorporates the contribution of adhesion energy between the layers of van der Waals heterostructures Yue2012 ; Wang2013a ; Dai2018 .
+More advanced models further integrate the equation of state of the encapsulated substance Zhilyaev2019 ; Iakovlev2019 ; Aslymov2020 ; Qu2023 and account for nonlinear effects Lyublinskaya2020 .
+These models have revealed that circular nanobubbles exhibit universal scaling behavior, characterized by a constant height-to-radius aspect ratio, $H/R$.
+Numerical simulations serve not only to validate analytical predictions but also to provide deeper insights into the physical properties of nanobubbles.
+For instance, it has been shown that spatial confinement and high internal pressure can induce the formation of internal crystalline structures, which may exhibit planar geometries Iakovlev2017 .
+Furthermore, the equilibrium shape of a nanobubble can critically depend on the nature of the encapsulated material Ghorbanfekr2017 .
+Notably, simulations by S.K.
+Jain et al.
+Jain2017 demonstrated that a nanobubble containing an ideal gas invariably assumes a circular shape with a constant aspect ratio of $H/R=0.204$.
+The capabilities of molecular dynamics (MD) simulations have enabled the exploration of more complex phase behaviors, such as the transition of confined argon into a liquid-crystalline state Korneva2023 .
+MD techniques have also been employed to model the mechanical indentation of a nanobubble by an atomic force microscope probe Faraji2022 .
+From a thermal management perspective, it has been demonstrated that the presence of nanobubbles can significantly reduce the interfacial thermal conductivity of layered van der Waals structures Qu2023a .
+Additionally, they have been identified as a potential cause of substantial thermal contraction in these materials Qu2023b .
+In this paper, all possible stationary states of graphene nanobubbles are numerically modeled.
+The paper is organized as follows.
+Section II describes the model employed.
 Section III presents the possible stationary states of the nanobubbles.
@@
 Consider a rectangular graphene sheet in which the chains of valent bonds along the $x$-axis have a ”zigzag” structure with $N_{x}$ nodes, and along the $y$-axis—an ”armchair” structure with $N_{y}$ nodes (the sheet lies parallel to the $z=0$ plane).
-Let us take a sheet with free edges—see Fig.
-1.
-The sheet consists of $N_{c}=N_{x}N_{y}/2-2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its length is $L_{x}=(N_{x}-1)r_{c}\sqrt{3}/2$ and its width is $L_{y}=(3N_{y}/4-1)r_{c}$ ($r_{c}=1.418$ Åis the C–C bond length).
-In all calculations, we will use a graphene sheet with $N_{x}=416$, $N_{y}=480$ (sheet dimensions: $50.963\times 50.906$ nm2, number of carbon atoms $N_{c}=99838$).
-We will assume that hydrogen atoms are attached to the edge carbon atoms, which participate in the formation of only two C–C bonds, forming united CH atoms with them of mass $M_{e}=13m_{p}$ (the mass of the internal sheet atoms is $M_{c}=12m_{p}$, where $m_{p}=1.66\times 10^{-27}$ kg is the proton mass).
+Consider a sheet with periodic and free boundary conditions (Fig.
+1).
+When periodic boundary conditions are applied, the sheet consists of $N_{c}=N_{x}N_{y}/2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its periodic lengths are $L_{x}=N_{x}r_{c}\sqrt{3}/2$ and $L_{y}=3N_{y}r_{c}/4$ ($r_{c}=1.418$ Å is the C–C bond length).
+When free boundary conditions are applied, the sheet consists of $N_{c}=N_{x}N_{y}/2-2$ carbon atoms and has a rectangular shape of size $L_{x}\times L_{y}$, where its length is $L_{x}=(N_{x}-1)r_{c}\sqrt{3}/2$ and its width is $L_{y}=(3N_{y}/4-1)r_{c}$.
+In all calculations, we use a graphene sheet with parameters $N_{x}=416$, $N_{y}=480$ (sheet dimensions: $51.086\times 51.048$ nm2, number of carbon atoms $N_{c}=99840$).
+For the graphene sheet with free boundaries, we assume that hydrogen atoms are attached to the edge carbon atoms, which participate in the formation of only two C–C bonds.
+These edge atoms, together with the attached hydrogen atoms, form united CH atoms with a mass of $M_{e}=13m_{p}$ (the mass of the interior sheet atoms is $M_{c}=12m_{p}$, where $m_{p}=1.66\times 10^{-27}$ kg is the proton mass).
+The shape of the nanobubble and its properties do not depend on the boundary conditions employed.
+We only note that the approach of a nanobubble to a free edge of the sheet can lead to its collapse (i.e., the release of gas atoms from under the sheet), an effect which is impossible under periodic boundary conditions.
 Figure 1: A rectangular graphene sheet with $N_{x}=10$, $N_{y}=12$.
@@
 where $\cos\varphi=({\bf u}_{n}-{\bf u}_{k},{\bf u}_{l}-{\bf u}_{k})/r_{nk}r_{kl}$, and $\varphi_{0}=2\pi/3$ is the equilibrium valent angle.
-Parameters $\alpha=1.7889$ nm-1 and $\epsilon_{\varphi}=1.3143$ eV can be found from the small amplitude oscillations spectrum of the graphene sheet Savin2008 .
+Parameters $\alpha=1.7889$ Å-1 and $\epsilon_{\varphi}=1.3143$ eV can be found from the small amplitude oscillations spectrum of the graphene sheet Savin2008 .
 Valence bonds between four adjacent carbon atoms $n$, $m$, $k$ , and $l$ constitute torsion angles, the potential energy of which can be defined as
@@
 where $\varepsilon_{0}$ is the interaction energy (adhesion energy) and $h_{0}$ is the equilibrium distance to the surface of the half-space.
-For a silicon dioxide (SiO2) substrate, the energy is $\varepsilon_{0}=0.074$ eV and the distance is $h_{0}=5$ Å  Koenig2011 .
-Table 1: Values of the parameters of the interaction potential with a flat substrate (5) for atoms C, Ar, Kr, Xe.
-C Ar Kr Xe
-$\varepsilon_{0}$ (eV)  0.074  0.1406  0.17242  0.19462
-$h_{0}$ (Å) 5.0 5.0 5.116 5.397
+For carbon atom and silicon dioxide (SiO2) substrate, the energy is $\varepsilon_{0}=0.074$ eV and the distance is $h_{0}=5$ Å  Koenig2011 .
+The values used for the remaining atoms are presented in Table 1; the parameters were obtained using the Lorentz-Berthelot mixing rules.
+Table 1: Values of the parameters of the interaction potential with a flat substrate (5) and masses for atoms C, He, Ne, Ar, Kr, Xe.
+C He Ne Ar Kr Xe
+$\varepsilon_{0}$ (eV)  0.074  0.04322  0.0797  0.1406  0.17242  0.19462
+$h_{0}$ (Å) 5.0 4.577 4.662 5.0 5.116 5.397
+$M_{g}$ ($m_{p}$) 12.0 4.0 20.18 39.95 83.80 131.29
 The Hamiltonian of the inert gas molecular subsystem has the form
 ${\cal H}_{2}=\sum_{n=N_{c}+1}^{N}\left[\frac{1}{2}M_{g}(\dot{\bf u}_{n},\dot{\bf u}_{n})+Z({\bf u}_{n})\right]+E_{2},$ (6)
-where $N=N_{c}+N_{g}$ is the total number of atoms in the two-component molecular system, the vectors $\{{\bf u}_{n}\}_{n=N_{c}+1}^{N}$ specify the coordinates of the gas atoms, and $M_{g}$ is the atomic mass ($M_{g}=39.96$, 83.80, 131.29 $m_{p}$ for Ar, Kr, and Xe atoms, respectively).
-The first term in the sum (6) represents the kinetic energy of the gas subsystem, the second term—the interaction energy of the gas atoms with the flat substrate (the values of the interaction potential parameters (5) for the gas atoms are given in Table 1).
+where $N=N_{c}+N_{g}$ is the total number of atoms in the two-component molecular system, the vectors $\{{\bf u}_{n}\}_{n=N_{c}+1}^{N}$ specify the coordinates of the gas atoms, and $M_{g}$ is the atomic mass (see Table 1).
+The first term in the sum (6) represents the kinetic energy of the gas subsystem, the second term—the interaction energy of the gas atoms with the flat substrate.
 The last term,
@@
 defines the interaction energy between the gas atoms, where $r_{nk}=|{\bf u}_{n}-{\bf u}_{k}|$ is the distance between atoms $n$ and $k$.
-The interaction between atoms is described by the Lennard-Jones (6,12) potential:
+The interaction between atoms is described by the Lennard-Jones (12-6) potential:
 $V(r)=4\epsilon[(\sigma/r)^{12}-(\sigma/r)^{6}].$ (8)
-The values of the interaction potential parameters for various pairs of atoms are given in Table 2.
-Table 2: Values of the interaction potential parameters (8) for pairs of atoms CAr, CKr, CXe, ArAr, KrKr, XeXe Iakovlev2017 ; Rappe1992 ; Vogt2014 .
-CAr CKr CXe ArAr KrKr XeXe
-$\varepsilon_{1}$ (eV)  0.00525  0.00643  0.00724  0.010  0.015  0.019
-$\sigma$ (Å) 3.3965 3.4965 3.7465 3.40 3.60 4.10
+The values of the interaction potential parameters for various pairs of atoms are given in Table 2. The LJ interactions were truncated at a cutoff distance of 30 Å.
+Table 2: Values of the interaction potential parameters (8) for various pairs of atoms Iakovlev2017 ; Rappe1992 ; Vogt2014 .
+CHe CNe CAr CKr CXe
+$\varepsilon_{1}$ (eV)  0.00161  0.00297  0.00525  0.00643  0.00724
+$\sigma$ (Å) 3.0165 3.0915 3.3965 3.4965 3.7465
+HeHe NeNe ArAr KrKr XeXe
+$\varepsilon_{1}$ (eV) 0.00094 0.0032 0.010 0.015 0.019
+$\sigma$ (Å) 2.64 2.79 3.40 3.60 4.10
 The interaction energy between the gas atoms and the graphene sheet is given by the sum
@@
 $E=\sum_{n=1}^{N_{c}}Q_{n}+\sum_{n=1}^{N}Z(\mathbf{u}_{n})+E_{2}+E_{3}.$ (11)
-Figure 2: Stationary $l$-layer states of encapsulated $N_{g}=2909$ argon atoms with $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
+Figure 2: Stationary $l$-layer states of encapsulated $N_{g}=4000$ argon atoms with $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
 For clarity of presentation, only one half of the graphene sheet is shown.
 The graphene sheet is depicted in gray.
-For argon atoms, their vertical displacement is indicated by color (blue corresponds to the minimum displacement with coordinate $z=h_{0}$, red corresponds to the maximum displacement with $z=h_{0}+10$ Å).
-Figure 3: Distribution of vertical displacements $p(z)$ for the stationary states of encapsulated $N_{g}=3936$ krypton atoms with number of layers $l=1$, …, 5.
+For argon atoms, their vertical displacement is indicated by color.
+Figure 3: Distribution of vertical displacements $p(z)$ for the stationary states of encapsulated $N_{g}=4000$ argon atoms with number of layers $l=1$, …, 5.
 More detailed discussion and motivation of our choice of the interaction potentials (2), (3), (4) can be found in Savin2010 .
@@
 By varying the initial configuration, all possible stable stationary states of the encapsulated molecular system can be obtained.
-Let $\{{\bf u}_{n}^{0}=(x_{n}^{0},y_{n}^{0},z_{n}^{0})\}_{n=1}^{N}$ be a solution to problem (12).
-The corresponding stationary state is characterized by its energy $E(\{{\bf u}_{n}^{0}\}_{n=1}^{N})$, the maximum value of the vertical displacements of the graphene sheet over the encapsulated group of atoms (the height of the graphene bulge)
+Let $\{{\bf u}_{n}^{0}=(x_{n}^{0},y_{n}^{0},z_{n}^{0})\}_{n=1}^{N}$ denote a solution to problem (12).
+The corresponding stationary state is characterized by its total energy $E(\{{\bf u}_{n}^{0}\}_{n=1}^{N})$, the maximum vertical displacement of the graphene sheet above the encapsulated group of atoms (i.e., the height of the graphene bulge)
 $H=\max_{n=1,\ldots,N_{c}}(z_{n}^{0}-h_{0}),$
 the radius $R$ of the bulge boundary, as well as the density $d$ and pressure $P$ inside the nanobubble.
-We will consider that a sheet atom participates in forming the upper surface of the nanobubble if its vertical displacement $\Delta z_{n}=z_{n}^{0}-h_{0}>h_{1}$, where the displacement threshold is $h_{1}=1.5$ Å.
+Figure 4: Stationary $l$-layer states of a nanobubble with $N_{g}=4000$ krypton atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
+A top view is shown.
+Color indicates the distribution of deformation energy across the graphene sheet.
+Figure 5: Stationary $l$-layer states of a nanobubble with $N_{g}=4000$ xenon atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
+A top view is shown.
+Color indicates the distribution of pressure across the graphene sheet.
+The average pressure values are $P=0.976$, 0.948, 0.977, 0.989, and 0.995 GPa.
+We consider a sheet atom to participate in forming the upper surface of the nanobubble if its vertical displacement $\Delta z_{n}=z_{n}^{0}-h_{0}>h_{1}$, where the displacement threshold $h_{1}$ is 0.97, 1.17, 1.5, 1.6, and 1.8 Å  for He, Ne, Ar, Kr, and Xe atoms, respectively.
 The center of the nanobubble then has coordinates
@@
 $N_{i}=\sum_{n=1\atop\Delta z_{n}>h_{1}}^{N_{c}}1.$
+Table 3: Dependence on the number of encapsulated helium atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
+The number of layers in the ground state is highlighted in bold.
+$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
+125 1 0 1.933 2.19 13.58 0.52 0.161
+250 1 0 1.641 2.12 19.86 0.55 0.107
+2 0.0076 1.474 3.82 19.09 0.50 0.200
+500 1 0 1.422 2.11 28.31 0.57 0.075
+2 0.0049 1.227 4.33 26.09 0.51 0.166
+1 0 1.252 2.10 40.54 0.59 0.052
+1000 2 0.0039 1.051 4.54 36.42 0.51 0.125
+3 0.0038 0.979 6.24 33.50 0.51 0.186
+1 0 1.152 2.16 57.55 0.60 0.038
+2 0.0026 0.930 4.45 50.40 0.52 0.088
+2000 3 0.0008 0.818 6.80 44.98 0.50 0.151
+1 0 1.070 2.17 92.48 0.58 0.023
+2 0.0005 0.897 4.49 72.01 0.55 0.062
+4000 3 -0.0014 0.758 6.80 65.84 0.51 0.103
+4 -0.0025 0.694 9.01 62.92 0.49 0.143
+5 -0.0031 0.652 10.44 60.11 0.49 0.174
+Table 4: Dependence on the number of encapsulated neon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
+$N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
+125 1 0 1.736 2.52 15.81 1.69 0.159
+250 1 0 1.450 2.51 22.89 1.80 0.110
+2 0.0101 1.476 4.64 21.03 1.71 0.221
+500 1 0 1.224 2.50 32.71 1.86 0.076
+2 0.0077 1.237 4.93 28.55 1.77 0.173
+1 0 1.098 2.51 46.06 1.94 0.054
+1000 2 0.0034 1.007 4.92 38.17 1.83 0.129
+3 0.0068 1.028 7.16 36.39 1.80 0.197
+1 0 1.001 2.50 65.37 1.97 0.038
+2 0.0022 0.839 4.93 52.07 1.87 0.095
+2000 3 0.0034 0.872 7.40 48.46 1.83 0.153
+4 0.0050 0.876 9.25 46.98 1.82 0.197
+1 0 0.932 2.50 92.48 2.01 0.027
+2 0.0011 0.727 4.93 72.01 1.90 0.068
+4000 3 0.0015 0.760 7.38 65.84 1.85 0.112
+4 0.0021 0.774 9.86 62.92 1.82 0.157
+5 0.0030 0.752 11.86 60.11 1.82 0.197
 It is convenient to define the base radius of the bubble as
 $R=\frac{1}{N_{r}}\sum_{n=1\atop\Delta z_{n}\in[h_{1},h_{2}]}^{N_{c}}r_{n},\quad N_{r}=\sum_{n=1\atop\Delta z_{n}\in[h_{1},h_{2}]}^{N_{c}}1.$
-Here, the second vertical displacement threshold is $h_{2}=2.5$ Å, and the distance to the center is $r_{n}=[(x_{n}^{0}-x_{c})^{2}+(y_{n}^{0}-y_{c})^{2}]^{1/2}$.
+Here, the distance to the center is
+$r_{n}=[(x_{n}^{0}-x_{c})^{2}+(y_{n}^{0}-y_{c})^{2}]^{1/2},$
+and the second vertical displacement threshold $h_{2}$ is 1.62, 1.95, 2.5, 2.7, and 3.1 Å  for He, Ne, Ar, Kr, and Xe atoms, respectively.
 The volume of the encapsulation cavity beneath the graphene sheet (the nanobubble volume) is
@@
 The density of gas atoms inside the nanobubble is $d=M_{g}N_{g}/V$.
-Figure 4: Stationary $l$-layer states of a nanobubble with $N_{g}=3936$ krypton atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
-A top view is shown.
-Color indicates the distribution of deformation energy across the graphene sheet (blue corresponds to zero deformation, red corresponds to deformation with an energy of 0.008 eV per atom).
-Figure 5: Stationary $l$-layer states of a nanobubble with $N_{g}=2909$ xenon atoms at $l=1$ (a), 2 (b), 3 (c), 4 (d), and 5 (e).
-A top view is shown.
-Color indicates the distribution of pressure across the graphene sheet (blue corresponds to zero pressure, red corresponds to a pressure of $p_{n}=5$ GPa).
-The average pressure values are $P=1.09$, 0.90, 0.86, 0.84, and 0.85 GPa.
 In the stationary state, a force acts on the $n$-th atom of the graphene sheet from the side of the encapsulated atoms:
-${\bf F}_{n}=\frac{4\epsilon_{0}}{\sigma^{2}}\sum_{k=N_{c}+1}^{N}\left[2(\sigma/r_{nk})^{14}-(\sigma/r_{nk})^{8}\right]({\bf u}_{n}^{0}-{\bf u}_{k}^{0}),$
+${\bf F}_{n}=\frac{24\epsilon_{0}}{\sigma^{2}}\sum_{k=N_{c}+1}^{N}\left[2(\sigma/r_{nk})^{14}-(\sigma/r_{nk})^{8}\right]({\bf u}_{n}^{0}-{\bf u}_{k}^{0}),$
 where the distance between atoms is $r_{nk}=|{\bf u}_{n}^{0}-{\bf u}_{k}^{0}|$.
@@
 Table 5: Dependence on the number of encapsulated argon atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
-Values for the ground states are highlighted in italics.
 $N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
-188 1 0 1.428 3.23 25.33 1.70 0.128
-2 0.0188 1.307 5.90 22.98 1.64 0.128
-260 1 0 1.257 3.23 29.84 1.76 0.108
-2 0.0137 1.302 6.12 26.33 1.68 0.232
-398 1 0 1.171 3.23 36.71 1.80 0.088
-2 0.0063 1.149 6.23 31.38 1.74 0.199
-614 1 0 1.080 3.23 45.71 1.85 0.071
-2 0.0015 1.030 6.22 37.64 1.79 0.165
-1 0 1.012 3.23 57.33 1.87 0.056
-962 2 -0.0018 0.950 6.22 46.57 1.82 0.165
-3 0.0010 0.953 9.11 44.30 1.79 0.206
-1 0 0.974 3.23 67.34 1.90 0.048
-1335 2 -0.0046 0.866 6.22 53.08 1.85 0.117
-3 -0.0046 0.892 9.24 50.68 1.80 0.183
-1 0 0.942 3.23 78.83 1.91 0.041
-1828 2 -0.0063 0.801 6.22 61.26 1.87 0.102
-3 -0.0048 0.793 9.23 56.41 1.83 0.164
-4 -0.0003 0.796 12.02 54.32 1.81 0.221
-1 0 0.906 3.24 99.46 1.93 0.033
-2 -0.0083 0.741 6.23 75.84 1.89 0.082
-2909 3 -0.0063 0.767 9.25 73.10 1.85 0.127
-4 -0.0065 0.708 12.20 66.67 1.83 0.183
-5 -0.0032 0.705 14.55 64.18 1.82 0.227
-1 0 0.885 3.24 115.70 1.94 0.028
-2 -0.0084 0.723 6.24 88.75 1.90 0.070
-3936 3 -0.0084 0.723 9.24 82.77 1.86 0.112
-4 -0.0097 0.696 12.24 76.09 1.84 0.161
-5 -0.0079 0.678 15.02 73.79 1.83 0.204
+125 1 0 1.601 3.24 20.65 1.64 0.157
+250 1 0 1.329 3.23 29.33 1.75 0.110
+2 0.0156 1.446 6.09 26.01 1.68 0.234
+500 1 0 1.159 3.23 41.22 1.82 0.078
+2 0.0045 1.231 6.22 34.60 1.77 0.180
+1 0 1.032 3.23 58.35 1.87 0.055
+1000 2 -0.0028 1.067 6.21 46.94 1.82 0.132
+3 0.0036 1.098 9.16 44.54 1.79 0.206
+1 0 0.946 3.23 82.38 1.91 0.039
+2 -0.0061 0.939 6.22 64.20 1.87 0.097
+2000 3 -0.0045 0.974 9.22 58.45 1.83 0.158
+4 -0.0006 1.007 12.04 56.71 1.81 0.212
+1 0 0.886 3.23 116.39 1.94 0.028
+2 -0.0089 0.845 6.22 88.35 1.90 0.070
+4000 3 -0.0098 0.899 9.22 80.29 1.86 0.115
+4 -0.0085 0.919 12.24 76.26 1.84 0.161
+5 -0.0060 0.920 15.09 72.65 1.83 0.208
 Table 6: Dependence on the number of encapsulated krypton atoms $N_{g}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble.
 $N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
-188 1 0 1.442 3.35 27.46 3.02 0.122
-2 0.0234 1.509 5.99 25.78 2.86 0.232
-260 1 0 1.357 3.35 32.23 3.11 0.084
-2 0.0179 1.389 6.45 28.48 2.96 0.226
-398 1 0 1.284 3.35 39.91 3.17 0.084
-2 0.0086 1.240 6.53 33.90 3.05 0.193
-1 0 1.194 3.36 49.24 3.26 0.068
-614 2 0.0041 1.133 6.52 40.75 3.13 0.160
-3 0.0145 1.151 9.07 39.78 3.06 0.228
-1 0 1.131 3.35 61.63 3.30 0.054
-962 2 -0.0006 1.063 6.53 49.85 3.19 0.131
-3 0.0068 1.023 9.61 47.47 3.12 0.202
-1 0 1.098 3.35 72.43 3.35 0.046
-1335 2 -0.0033 0.982 6.53 57.02 3.24 0.115
-3 0.0015 0.964 9.72 53.34 3.16 0.182
-4 0.0930 0.959 12.49 51.15 3.14 0.206
-1 0 1.069 3.35 84.72 3.37 0.040
-1828 2 -0.0055 0.944 6.54 66.20 3.26 0.099
-3 -0.0029 0.892 9.70 60.17 3.20 0.161
-4 0.0005 0.941 12.41 60.66 3.17 0.205
-1 0 1.038 3.35 106.65 3.41 0.031
-2 -0.0080 0.872 6.54 81.66 3.31 0.080
-2909 3 -0.0070 0.852 9.71 74.79 3.23 0.130
-4 -0.0047 0.813 12.94 71.05 3.19 0.182
-5 -0.0015 0.877 15.41 72.16 3.18 0.214
-1 0 1.018 3.36 123.73 3.44 0.027
-2 -0.0096 0.827 6.55 93.70 3.34 0.070
-3936 3 -0.0094 0.806 9.71 85.27 3.25 0.114
-4 -0.0082 0.761 12.91 79.40 3.21 0.163
-5 -0.0061 0.796 15.98 78.45 3.20 0.204
-The solution of problem (12) showed that each stationary state of a nanobubble always has a strictly defined number of atomic layers, $l$, in the cluster of internal atoms—see Figs.
+125 1 0 1.599 3.48 22.31 2.80 0.156
+2 0.0296 1.714 5.72 21.28 2.67 0.269
+250 1 0 1.368 3.48 31.44 2.97 0.182
+2 0.0179 1.480 6.57 27.61 2.87 0.238
+500 1 0 1.200 3.48 44.27 3.10 0.079
+2 0.0043 1.293 6.64 37.07 3.01 0.179
+1 0 1.099 3.48 62.51 3.18 0.056
+1000 2 -0.0052 1.148 6.64 50.02 3.12 0.133
+3 0.0032 1.162 9.79 46.38 3.08 0.211
+1 0 1.021 3.48 88.31 3.25 0.039
+2000 2 -0.0095 1.035 6.64 68.59 3.20 0.097
+3 -0.0073 1.068 9.82 62.33 3.15 0.158
+4 0.0024 1.085 12.88 60.19 3.12 0.214
+1 0 0.968 3.48 124.55 3.30 0.028
+2 -0.0128 0.951 6.64 94.53 3.26 0.070
+4000 3 -0.0134 0.997 9.82 85.73 3.20 0.115
+4 -0.0116 1.019 13.02 81.56 3.17 0.160
+5 -0.0080 1.025 16.09 77.39 3.16 0.208
+Solving problem (12) reveals that each stationary state of a nanobubble possesses a strictly defined number of atomic layers, $l$, within the internal cluster—see Figs.
 2 and 3.
-For a number of encapsulated atoms $N_{g}<4000$, stationary states with $l=1,\ldots,6$ layers are possible.
-Each internal atom lies within one of the layers (see Fig.
+For systems with up to $N_{g}\leq 4000$ encapsulated atoms, stationary states with l=1,,6l=1,...,6 layers are possible.
+Each internal atom resides within one of these layers (Fig.
 3).
-The layers have a flat, circular shape (ranging from the shape of a regular hexagon to a circle).
-The layers are concentrically stacked on top of each other and form a symmetric $l$-stepped pyramid with a flat top (the plane of the layers is always parallel to the substrate plane).
-Atoms within each layer are in the same plane, with only edge atoms displaced downwards, smoothing the stepped shape of the pyramid.
-Atoms Ar, Kr, and Xe differ in size, hence the interlayer distances for them are $\Delta h=3.0$, 3.2, and 3.6 Å, respectively.
-For argon atoms, two- and three-layer structures at $N_{g}=2463$ were first obtained in Iakovlev2017 .
+The layers are flat and circular in shape.
+They are concentrically stacked, forming a symmetric $l$-stepped pyramid with a flat top, and remain parallel to the substrate plane.
+While atoms within each layer lie in the same plane, the edge atoms are displaced downward, effectively smoothing the stepped profile of the pyramid.
+The size of the inert gas atoms varies across the series, resulting in distinct interlayer distances: $\Delta h=2.2$, 2.5, 3.0, 3.2, and 3.6 Å for He, Ne, Ar, Kr, and Xe, respectively.
+For argon atoms specifically, two- and three-layer structures at $N_{g}=2463$ were first reported in Iakovlev2017 .
 Covering the $l$-stepped pyramid from above with a graphene sheet is achieved through its local stretching.
@@
 The most significant stretching of C–C bonds occurs over the layer boundaries, where the deformation energy can reach values of 0.01 eV per carbon atom.
+Here, valence bonds can elongate by 1%.
+This is in agreement with experimental measurements on graphene bubbles Georgiou2011 ; Zabel2011 .
 As seen in Fig.
 4, the zones of sheet stretching strictly correspond to the stepped structure of the pyramid of encapsulated atoms.
-For example, for $N_{g}=3936$ and $l$-layer states with $l=1,2,3$, stretching of the sheet occurs only over the edges of the layers.
+For example, for $N_{g}=4000$ and $l$-layer states with $l=1,2,3$, stretching of the sheet mainly occurs above the edges of the layers.
 For states with a larger number of layers ($l=4,5$), stretching occurs more uniformly across the entire surface of the nanobubble, but the maximum values are always reached at its base (over the edge of the first layer).
@@
 $N_{g}$   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
-188 1 0 1.476 3.57 31.54 3.52 0.113
-2 0.0278 1.492 6.87 28.87 3.28 0.231
-260 1 0 1.361 3.58 36.85 3.60 0.097
-2 0.0216 1.367 7.15 32.62 3.34 0.219
-398 1 0 1.296 3.58 45.58 3.66 0.079
-2 0.0061 1.096 7.21 38.37 3.45 0.188
-1 0 1.227 3.59 56.15 3.75 0.064
-614 2 -0.0009 1.133 7.23 45.86 3.50 0.158
-3 0.0108 1.165 10.36 44.91 3.41 0.213
-1 0 1.180 3.58 70.11 3.80 0.051
-962 2 -0.0068 1.083 7.23 56.27 3.55 0.128
-3 -0.0013 1.015 10.76 53.26 3.45 0.202
-1 0 1.146 3.58 82.44 3.84 0.043
-1335 2 -0.0117 1.006 7.24 65.08 3.57 0.111
-3 -0.0104 0.965 10.88 60.08 3.49 0.181
-4 00.0012 0.980 14.04 58.02 3.45 0.242
-1 0 1.125 3.59 96.27 3.86 0.037
-1828 2 -0.0160 0.953 7.25 74.47 3.62 0.097
-3 -0.0306 0.880 10.93 67.94 3.51 0.161
-4 -0.0124 0.925 14.36 65.59 3.48 0.219
-1 0 1.090 3.60 120.99 3.90 0.030
-2 -0.0198 0.897 7.27 92.58 3.65 0.079
-2909 3 -0.0238 0.860 10.89 83.19 3.55 0.131
-4 -0.0228 0.841 14.55 79.27 3.50 0.184
-5 -0.0171 0.846 17.81 76.76 3.47 0.232
-1 0 1.077 3.60 140.68 3.91 0.026
-2 -0.0218 0.872 7.26 106.36 3.67 0.068
-3936 3 -0.0272 0.828 10.88 95.06 3.56 0.114
-4 -0.0286 0.791 14.53 87.11 3.52 0.167
-5 -0.0255 0.794 17.79 88.19 3.47 0.202
-6 -0.0233 0.804 20.81 85.72 3.48 0.243
-The maximum possible number of layers, $l_{m}$, depends on the number of internal atoms, $N_{g}$ ($l_{m}$ increases with increasing $N_{g}$).
-For argon atoms, only single-layer packings ($l_{m}=1$) are possible for $N_{g}\leq 150$, while for $150<N_{g}<900$, single- and double-layer packings ($l_{m}=2$) are possible.
-For $900<N_{g}<1800$, the maximum number of layers is $l_{m}=3$; for $1800<N_{g}<2800$, it is $l_{m}=4$; and for $2800<N_{g}<4000$, it is $l_{m}=5$.
+125 1 0 1.545 4.01 25.39 2.99 0.158
+2 0.0347 1.646 6.69 24.09 2.85 0.278
+250 1 0 1.324 4.01 35.82 3.14 0.112
+2 0.0134 1.411 7.54 31.48 3.04 0.240
+500 1 0 1.182 4.02 50.36 3.27 0.080
+2 -0.0081 1.238 7.64 41.87 3.19 0.182
+3 0.0087 1.265 10.51 40.62 3.13 0.259
+1 0 1.087 4.02 71.04 3.36 0.057
+1000 2 -0.0226 1.113 7.64 56.73 3.30 0.135
+3 -0.0154 1.141 11.23 52.79 3.25 0.213
+1 0 1.021 4.02 100.34 3.42 0.040
+2000 2 -0.0298 1.012 7.64 77.41 3.38 0.099
+3 -0.0322 1.047 11.27 70.32 3.34 0.160
+4 -0.0269 1.060 14.74 68.33 3.30 0.216
+1 0 0.976 4.02 141.59 3.48 0.028
+2 -0.0350 0.948 7.65 107.08 3.44 0.071
+4000 3 -0.0411 0.977 11.28 96.85 3.38 0.116
+4 -0.0412 0.989 14.94 91.93 3.35 0.163
+5 -0.0380 0.995 18.43 87.79 3.34 0.210
+6 -0.0345 0.996 20.81 86.63 3.33 0.240
+Table 8: Dependence on the adhesion energy $\varepsilon_{0}$ and the number of layers $l$ of the relative specific energy $\Delta E$, pressure $P$, height $H$, radius $R$, and density $d$ of the stationary state of the nanobubble with $N_{g}=4000$ argon atoms.
+$\varepsilon_{0}$ (eV)   $l$ $\Delta E$ (eV) $P$ (GPa) $H$ (Å) $R$ (Å) $d$ (g/cm3) $H/R$
+1 0 0.684 3.27 119.76 1.81 0.027
+2 0.0165 0.709 6.27 92.22 1.78 0.068
+0.037 3 0.0035 0.729 9.27 83.28 1.76 0.111
+4 0.0061 0.728 12.30 79.03 1.75 0.156
+5 0.0095 0.727 15.00 75.83 1.74 0.198
+1 0 1.407 3.17 112.57 2.12 0.028
+2 -0.0247 1.118 6.14 85.36 2.05 0.072
+0.148 3 -0.0302 1.209 9.14 77.63 2.00 0.118
+4 -0.0317 1.241 12.18 73.43 1.96 0.166
+5 -0.0310 1.249 15.08 69.62 1.95 0.217
+The maximum possible number of layers, $l_{m}$, depends on the number of encapsulated atoms, $N_{g}$, increasing monotonically with $N_{g}$.
+For helium atoms, the following behavior is observed: for $N_{g}<N_{1}=188$, only single-layer packings are possible ($l_{m}=1$); for $N_{1}<N_{g}<N_{2}=850$, both single- and double-layer packings are possible ($l_{m}=2$); for $N_{2}<N_{g}<N_{3}=2500$, the maximum number of layers is $l_{m}=3$; for $N_{3}<N_{g}<N_{4}=3500$, it is $l_{m}=4$; and for $N_{4}<N_{g}<4000$, it is $l_{m}=5$.
+The boundary values of the number of atoms, $N_{k}$ for the other noble gas are as follows: $N_{k}=182$, 650, 1500, 3500 for Ne; $N_{k}=145$, 550, 1600, 2500 for Ar; $N_{k}=123$, 450, 1300, 2400 for Kr; and $N_{k}=113$, 430, 1200, 2200 for Xe.
 A single-layer packing exists for any number of atoms.
 Therefore, the relative energy of a stationary state can be measured from the energy of the single-layer packing: $\Delta E=(E_{l}-E_{1})/N_{g}$, where $E_{l}$ is the energy of the $l$-layer stationary state.
-The dependence of the relative specific energy $\Delta E$, internal pressure $P$, height $H$, radius $R$, and density $d$ of the nanobubble’s stationary state on the number of argon atoms $N_{g}$ is given in Table 3.
-As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$.
-For $900<N_{g}<3000$, two-layer states become more favorable, and for $N_{g}>3000$, three- and four-layer states are more favorable.
+The dependence of the relative specific energy $\Delta E$, internal pressure $P$, height $H$, radius $R$, and density $d$ of the nanobubble’s stationary state on the number of helium atoms $N_{g}$ is given in Table 3.
+As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}\leq 2000$.
 The pressure inside the encapsulation monotonically decreases with an increasing number of atoms.
 With an increasing number of layers, a slight decrease in pressure and density is observed.
-The height of the nanobubble, $H$, always takes a discrete set of values proportional to the number of layers $l$: $H\approx h_{0}+(l-1)\Delta h$.
-For krypton atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of internal atoms $N_{g}$ is presented in Table 4.
-Here, for $N_{g}\leq 140$, the maximum number of layers is $l_{m}=1$; for $140<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$.
-As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<900$, while two-layer states are favorable for $900<N_{g}<4000$.
-Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms.
+The height of the nanobubble, $H$, always takes a discrete set of values proportional to the number of layers $l$: $H\approx l\Delta h$.
+For neon, argon, krypton and xenon atoms, the dependencies of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of encapsulated atoms $N_{g}$ are presented in Tables 4, 5, 6 and 7, respectively.
+As can be seen from the tables, for all encapsulated atoms, the pressure inside the nanobubble decreases monotonically with an increasing number of atoms.
 Pressure and density reach their maximum values for single-layer states.
-For xenon atoms, the dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on $N_{g}$ is presented in Table 5.
-Here, for $N_{g}\leq 116$, the maximum number of layers is $l_{m}=1$; for $116<N_{g}<600$, $l_{m}=2$; for $600<N_{g}<1300$, $l_{m}=3$; for $1300<N_{g}<2800$, $l_{m}=4$; for $2800<N_{g}<3800$, $l_{m}=5$; and for $3800<N_{g}<4000$, $l_{m}=6$.
-As can be seen from the table, single-layer stationary states are the most energetically favorable for $N_{g}<600$, two-layer states are favorable for $600<N_{g}<1400$, and three-layer states are favorable thereafter.
-Here too, the pressure inside the nanobubble monotonically decreases with an increasing number of atoms, and pressure and density reach their maximum values for single-layer packings.
-Figure 6: Dependence on temperature $T$ of (a) height $H$, (b) density $d$, and (c) internal pressure $P$ for a graphene nanobubble with $N_{g}=2909$ xenon atoms.
+For neon atoms, single-layer stationary states are the most energetically favorable across the entire range investigated ($N_{g}\leq 4000$).
+For argon atoms, single-layer stationary states are the most energetically favorable only for $N_{g}<800$.
+For $800<N_{g}\leq 3000$, two-layer states become more favorable, and for $N_{g}>3000$, three-layer states are preferred.
+For krypton atoms, single-layer stationary states are the most energetically favorable for $N_{g}<700$.
+Two-layer states are favorable for $700<N_{g}<3000$, and three-layer states are favored for $N_{g}>3000$.
+For xenon atoms, single-layer stationary states are the most energetically favorable for $N_{g}<500$, while two-layer states are favorable for $500<N_{g}<2000$.
+For $N_{g}>2000$, three- and four-layer states become increasingly favorable.
+Figure 6: Dependence on temperature $T$ of (a) height $H$, (b) density $d$, and (c) internal pressure $P$ for a graphene nanobubble with $N_{g}=4000$ argon atoms.
 Curves 1, 6, 11 correspond to the state with $l=1$; curves 2, 7, 12 – to $l=2$, …, and curves 5, 10, 15 – to $l=5$ layers.
-The dashed vertical lines correspond to the transition temperatures between states $l\rightarrow 4$: $T_{1}=960$, $T_{2}=390$, $T_{3}=330$, and $T_{5}=210$ K.
-The dotted vertical lines correspond to the melting temperature of the single-layer (420 K) and four-layer state (270 K).
-Figure 7: Dependence on temperature $T$ of the vertical displacement distribution $p(z)$ for $l$-layer states of $N_{g}=2909$ internal xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
+The dashed vertical lines correspond to the transition temperatures between states $l\rightarrow 4$: $T_{1}=630$, $T_{2}=270$, $T_{3}=180$, and $T_{5}=120$ K.
+The dotted vertical lines correspond to the melting temperature of the single-layer ($T_{1,m}=255$ K) and four-layer state ($T_{4,m}=160$ K).
+Figure 7: Dependence on temperature $T$ of height of graphene nanobubble $H$ with $N_{g}=4000$ internal (a) helium, (b) neon, (c) krypton and (d) xenon atoms (curves 1,…, 5 correspond to the state with l=1l=1,…, 5 layers.
+Figure 8: Dependence on temperature $T$ of the vertical displacement distribution $p(z)$ for $l$-layer states of $N_{g}=4000$ internal argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
+To understand how the shape of a nanobubble’s stationary state depends on the value of the interaction energy between the graphene sheet and the substrate, we also consider states with twice weaker and twice stronger interactions: at $\epsilon_{0}=0.037$ and $0.148$ eV.
+The weak interaction is characteristic of substrates with low density.
+For example, for an Ih ice substrate, the energy is $\epsilon_{0}=0.029$ eV Savin2019 .
+The strongest interaction occurs with the surface of a nickel crystal [with Ni(111) surface], where $\varepsilon_{0}=0.133$ eV Lahiri2011 .
+Thus, the three values of adhesion energy $\varepsilon=0.037$, 0.074, and 0.148 eV correspond to weak, medium (the most typical), and strong interaction of the graphene sheet with the substrate.
+The numerical solution of the problem (12) showed that for all three characteristic values of the adhesion energy, a nanobubble containing $N_{g}=4000$ argon atoms can only exist in stationary states with l=1,,5l=1,...,5 layers; i.e., the maximum possible number of layers is $l_{m}=5$.
+The dependence of $\Delta E$, $P$, $H$, $R$, and $d$ on the number of layers is presented in Tables 5 and 8.
+As can be seen from the tables, the shape of the nanobubble’s stationary state depends only weakly on the value of $\epsilon_{0}$.
+A twofold increase in the adhesion energy leads to an increase in the internal pressure $P$, accompanied by a slight decrease in the nanobubble radius $R$.
+With weak adhesion, the most energetically favorable state is a single-layer nanobubble; with medium adhesion, a three-layer state; and with strong adhesion, a five-layer stationary state is the most favorable.
+Thus, an increase in the interaction energy between the graphene sheet and the substrate does not significantly change the step-like shape of the stationary states, but it noticeably increases the internal pressure and makes multilayer states energetically more favorable.
 IV Stability of Stationary States to Thermal Fluctuations
@@
 To verify their stability against thermal fluctuations, we performed simulations of nanobubble dynamics at temperatures $T\leq 1000$ K.
-Figure 8: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner xenon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
-Figure 9: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=2909$ inner argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
+Figure 9: Temperature dependence $T$ of the transverse displacement distribution $p(r)$ from the center of mass for $l$-layer states of $N_{g}=4000$ inner argon atoms: (a) $l=1$; (b) $l=2$; (c) $l=3$; and (d) $l=4$.
+Let us consider the dynamics of a molecular system in which a graphene sheet interacts with a Langevin thermostat.
 Dynamics of a thermalized graphene sheet with encapsulated atoms is described by the system of Langevin equations:
-$M_{n}\ddot{\bf u}_{n}=-\frac{\partial H}{\partial{\bf u}_{n}}-\Gamma M_{n}\dot{\bf u}_{n}+\Xi_{n},~~n=1,...,N,$ (13)
+$M_{n}\ddot{\bf u}_{n}$ $=$ $-\frac{\partial H}{\partial{\bf u}_{n}}-\Gamma M_{n}\dot{\bf u}_{n}+\Xi_{n},~~n=1,...,N_{c},$~~~ (13)
+$M_{n}\ddot{\bf u}_{n}$ $=$ $-\frac{\partial H}{\partial{\bf u}_{n}},~~n=N_{c}+1,...,N,$ (14)
 where the mass $M_{n}=M_{c}$ for $n=1,...,N_{c}$ and $M_{n}=M_{g}$ for $n=N_{c}+1,...,N$ (the total number of atoms is $N=N_{c}+N_{g}$).
@@
 This allows us to significantly reduce the numerical integration time required to reach an equilibrium state and obtain reliable averages.
-We numerically integrate the system of equations of motion (13) with initial conditions corresponding to a stationary nanobubble state:
+We numerically integrate the system of equations of motion (13), (14) with initial conditions corresponding to a stationary nanobubble state:
 $\{{\bf u}_{n}(0)={\bf u}_{n}^{0},~\dot{\bf u}_{n}(0)=0\}_{n=1}^{N}.$ (15)
 The equations of motion Eq.
-(13) with initial conditions (14) are solved numerically using the velocity Verlet method Verlet1967 .
+(13), (14) with initial conditions (15) are solved numerically using the velocity Verlet method Verlet1967 .
 A time step of 1 fs is used in the simulations, as further reduction of the time step has no appreciable effect on the results.
-Numerical simulations of the dynamics showed that among the $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases.
-All other states transition into this ground state at a specific temperature.
+Numerical simulations of the dynamics showed that among the $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid configuration as the temperature increases.
+All other states transition into this ground state at a characteristic temperature $T_{l}$.
 The dependence of the nanobubble’s height $H$, density $d$, and internal pressure $P$ on temperature is shown in Fig.
 6.
+As can be seen from the figure, for a bubble with $N_{g}=4000$ argon atoms, a single-layer packing ($l=1$) persists up to a very high temperature $T_{1}=630$ K.
+Further temperature increase leads to its transition into a layerless molten configuration, accompanied by a sharp increase in height $H$, a decrease in density $d$, and internal pressure $P$.
+Two-layer atom packings ($l=2$) persist up to a temperature $T_{2}=270$ K, at which they transition into a four-layer packing.
+This transition is also accompanied by a sharp increase in height and a decrease in density.
+Three-layer packings ($l=3$) persist up to a temperature of $T_{3}=180$ K, and five-layer packings ($l=5$) persist up to $T_{5}=120$ K (see Fig.
+6 (a)).
+Here, the ground state is the four-layer $(l=4)$ packing.
+Note that with weak adhesion $\varepsilon_{0}=0.037$ eV, the ground state is the three-layer ($l=3$) state, while with strong adhesion $\varepsilon_{0}=0.148$ eV, it is the five-layer ($l=5$) state.
 […]
-In Tables 3, 4, 5, the rows corresponding to the ground states are given in italics.
+In Tables 3 to 8, the number of layers $l$ corresponding to the ground states is highlighted in bold.
 Note that the ground states are not always the most energetically favorable.
 Their key feature is the smooth (non-abrupt) change in their shape with increasing temperature.
+The ground state can also be obtained from the layerless configuration of the nanobubble at $T=900$ K by gradually reducing the thermostat temperature to zero.
 Conformational changes of the encapsulated atom cluster are reflected in the change in the distribution of atoms by their vertical $\{z_{n}\}_{n=N_{c}}^{N}$ (Fig.
-7) and radial (in-plane) displacements from the center of mass:
+8) and radial (in-plane) displacements from the center of mass:
 $\{r_{n}=[(x_{n}-x_{c})^{2}+(y_{n}-y_{c})^{2}]^{1/2}\}_{n=N_{c}+1}^{N}$
-(Figs. 8, 9).
+(Figs. 9).
 As can be seen from Fig.
@@
 Radial displacements also show changes in the distribution shape at the transition temperature $T_{l}$.
-The transition to a state with a larger number of layers leads to a sharp decrease in the diameter of the inner atom cluster—see Figs.
-8 (a, b, c) and 9 (a, b, c).
-For the ground state with $l=4$ layers, only a continuous, gradual increase in the cluster diameter (its thermal expansion) occurs—see Figs.
-8 (d) and 9 (d).
+The transition to a state with a larger number of layers leads to a sharp decrease in the diameter of the inner atom cluster—see Fig.
+9 (a,b,c).
+For the ground state with $l=4$ layers, only a continuous, gradual increase in the cluster diameter (its thermal expansion) occurs—see Fig.
+9 (d).
 The shape of the radial distribution also allows one to judge the phase state of the cluster.
@@
 Upon reaching the melting temperature $T_{l,m}$, the discrete distribution shape is replaced by a continuous one, characteristic of a liquid state.
-For $N_{g}=2909$ xenon atoms, the melting of the single-layer cluster occurs at $T_{1,m}=420\div 450$ K, and of the four-layer cluster at $T_{4,m}=270\div 300$ K—see Fig.
-8.
-For argon atoms, the melting temperatures are $T_{1,m}=240\div 270$ K, $T_{4,m}=150\div 180$ K (see Fig.
-9), and for krypton $T_{1,m}=300\div 330$ K, $T_{4,m}=210\div 240$ K.
-Such high melting temperatures are explained by the high value of internal pressure—see Fig.
-6 (c).
-Figure 10: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble: with $N_{g}=962$ and 2909 argon atoms (curves 1 and 2); with $N_{g}=3936$ krypton atoms (curve 3); with $N_{g}=2909$ xenon atoms (curve 4).
-The existence of several stable states with different numbers of atom layers $l$ at low temperatures indicates that the nanobubbles here cannot have a universal (scalable) shape, and the height-to-radius ratio is not constant.
-Indeed, for a single-layer state ($l=1$), increasing the number of internal atoms $N_{g}$ does not change the nanobubble’s height ($H\approx 3.3$ Å, while its radius $R$ grows as $\sqrt{N_{g}}$; therefore, the ratio $H/R\rightarrow 0$ as $N_{g}\rightarrow\infty$.
-Based on the data in Tables 3, 4, and 5, the ratio $H/R$ can vary from 0 to 0.24 here.
-The violation of the scaling rule for graphene bubbles with a radius of $\sim 1$ nm was noted in the experimental study Villarreal2021 , where they were created via the implantation of noble gas ions He, Ne, and Ar.
+For $N_{g}=4000$ argon atoms, the melting of the single-layer cluster occurs at $T_{1,m}\in$(240,270), and of the four-layer cluster at $T_{4,m}\in(150,180)$ K—see Fig.
+9.
+For helium atoms, the melting temperatures are $T_{1,m}\approx 80$, $T_{4,m}\approx 48$, for neon $T_{1,m}\approx 135$, $T_{4,m}\approx 75$, for krypton $T_{1,m}\approx 315$, $T_{4,m}\approx 240$, and for xenon $T_{1,m}\approx 435$, $T_{4,m}\approx 285$ K.
+Figure 10: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble with (a) $N_{g}=500$, 1000, 2000, 4000 argon atoms (curves 1, 2, 3, 4) and (b) with $N_{g}=4000$ helium, neon, krypton, xenon atoms (curve 5, 6, 7, 8).
+Figure 11: Temperature dependence $T$ of the ratio $H/R$ for the ground state of the nanobubble with $N_{g}=4000$ argon atoms for adhesion $\epsilon_{0}=0.037$, 0.074, 0.148 eV (curves 1, 2, 3)
+The obtained values of the one-layer stability temperature and the ground state melting temperature do not depend on the number of atoms in the encapsulated cluster; the same values are obtained for $N_{g}=500$, 1000, 2000, and 3000.
+Therefore, it can be expected that the obtained estimates will also hold for larger clusters (for $N_{g}>4000$).
+Noble gases transition from the solid to the liquid state at very low temperatures (24, 84, 114, and 161 K for Ne, Ar, Kr, and Xe, respectively).
+The high melting temperatures of their encapsulated clusters are explained by the presence of high internal pressure and the effect of strong spatial confinement Faraji2022 .
+The performed simulations show that the encapsulated noble gas cluster (graphene nanobubble) ceases to be a multistable system only at high temperatures $T>T_{1}$.
+Thus, at room temperature ($T=300$ K), graphene nanobubbles of Ne, Ar, Kr, and Xe, in addition to the ground three-dimensional state, also have a stable quasi-two-dimensional one-layer state—see Figs.
+6 (a) and 7 (b, c, d).
+The existence of several stable states with different numbers of atomic layers $l$ at low temperatures indicates that nanobubbles in this regime cannot possess a universal (scalable) shape; consequently, the height-to-radius ratio is not constant.
+Indeed, for a single-layer state ($l=1$), increasing the number of encapsulated atoms $N_{g}$ leaves the nanobubble’s height unchanged ($H=2.2$, 2.5, 3.2, 3.5, and 4.0 Å for He, Ne, Ar, Kr, and Xe, respectively), while its radius $R$ grows as $\sqrt{N_{g}}$.
+The ratio $H/R$ therefore approaches zero as $N_{g}\rightarrow\infty$.
+Based on the data in Tables 3 to 8, the ratio $H/R$ can vary from 0 to 0.28 here.
+The violation of the scaling rule for graphene bubbles with a radius of $\sim 1$ nm was noted in the experimental study Villarreal2021 , where they were created via the implantation of noble gas ions (He, Ne, and Ar).
 As can be seen from the table data, the relation $H/R\approx 0.2$ holds only for the ground states of the nanobubble.
@@
 10.
-As can be seen from the figure, the value of $H/R$ increases monotonically with temperature, while the spread of values decreases: $H/R$ varies within the intervals [0.163, 0.206] at $T=0$ K; [0.192, 0.208] at $T=300$ K; [0.211, 0.217] at $T=600$ K.
-These estimates are in good agreement with the universal value $H/R=0.204$ obtained in Jain2017 for a graphene nanobubble with an ideal gas.
-Note that the study Korneva2023 reported an estimate of $H/R\approx 0.13$ for graphene nanobubbles with argon atoms at $T=2$ K, indicating that the simulation used not a ground stationary state of the nanobubble, but rather a state with a smaller number of layers.
+As can be seen from the figure, the value of $H/R$ increases monotonically with temperature, while the spread of values decreases: $H/R$ varies within the intervals [0.151, 0.206] at $T=0$; [0.174, 0.212] at $T=300$; [0.204, 0.220] at $T=600$; and [0.212, 0.222] at $T=900$ K.
+These estimates are in good agreement with the universal value $H/R=0.204$ obtained in Jain2017 for a graphene nanobubble containing an ideal gas.
+Note that the study Korneva2023 reported an estimate of $H/R\approx 0.13$ for graphene nanobubbles with argon atoms at $T=2$ K, indicating that the simulation used not the ground stationary state of the nanobubble, but rather a state with a smaller number of layers.
+Note that the shape of the nanobubble’s ground state depends on the interaction energy $\epsilon_{0}$ between the graphene sheet and the substrate.
+The temperature dependence of the $H/R$ ratio for three characteristic values of the adhesion energy $\epsilon_{0}=0.037$, $0.074$, $0.148$ eV is shown in Fig.
+11.
+As can be seen from the figure, an increase in the adhesion energy leads to a monotonic increase in the $H/R$ ratio.
+For example, at a temperature of $T=300$ K, the $H/R$ ratio can increase from $0.163$ for weak adhesion to $0.222$ for strong adhesion; at $T=600$ K — from $0.177$ to $0.238$; and at $T=900$ K — from $0.182$ to $0.247$.
+This monotonic dependence of the $H/R$ ratio on $\epsilon_{0}$ allows for experimental estimates of the adhesion between the graphene sheet and the substrate Wang2016 .
 V Conclusion
-Using the example of inert gas atoms (Ar, Kr, Xe), it has been shown that graphene nanobubbles located on flat substrates are multistable systems.
-A nanobubble can have many stable stationary states, each characterized by the number of layers $l$ in the internal atom cluster.
-Inside the nanobubble, each atom lies within one of the layers.
-The layers have a circular, flat shape, are concentrically stacked on top of each other, and form an $l$-stepped pyramid with a flat top.
-Atoms within each layer lie in a single plane, with only the edge atoms displaced downwards, smoothing the stepped shape of the pyramid.
-The covering of the pyramid from above by a graphene sheet occurs due to its local stretching.
-The sheet’s valence bonds are stretched only above the group of encapsulated atoms; outside this coverage zone, the sheet remains undeformed and tightly adheres to the substrate.
-The strongest stretching (up to one percent) occurs above the edges of the layers.
-The maximum possible number of layers $l_{m}$ increases monotonically with the number of internal atoms $N_{g}$.
-For argon atoms, only single-layer stationary states ($l_{m}=1$) are possible for $N_{g}\leq 150$; for $150<N_{g}<900$—single-layer and double-layer states ($l_{m}=2$); for $900<N_{g}<1800$, $l_{m}=3$; for $1800<N_{g}<2800$, $l_{m}=4$; and for $2800<N_{g}<4000$, $l_{m}=5$.
-The graphene sheet, due to van der Waals interaction with the substrate, presses the cluster of encapsulated atoms to substrate, creating pressures $P\sim 1$ GPa.
-The highest pressure is experienced by single-layer packings of internal atoms.
-The magnitude of the pressure decreases as the number of layers increases.
-Thermal fluctuation modeling shows that among all $l$-layer states of the nanobubble, there is always one ”ground” state, which smoothly transitions into a layerless liquid state as the temperature increases.
-The other states transition into this state at a specific temperature $T_{l}$.
-For example, for $N_{g}=2909$ encapsulated atoms, the ground state is the four-layer packing ($l=4$).
-For Xe atoms, the $l$-layer states persist up to temperatures $T_{l}=960$, 390, 330, 210 K; for Ar atoms, $T_{l}=600$, 270, 180, 30 K; and for Kr atoms—750, 360, 240, 150 K, where $l=1$, 2, 3, 5.
-The existence of several stable states with different numbers of layers in the internal atom cluster at low temperatures indicates that universal scaling does not hold for these nanobubbles.
-The height-to-radius ratio $H/R$ of the nanobubble is not constant and can vary from 0 to 0.24.
-The universal relation $H/R\approx 0.2$ holds only for the ground states.
+Using the example of inert gas atoms (He, Ne, Ar, Kr, Xe), it has been shown that graphene nanobubbles on flat substrates exhibit multistability.
+These bubbles can adopt multiple stable stationary states, each defined by a distinct number of layers, $l$, within an internal cluster of atoms.
+Within the nanobubble, each atom is confined to one of these layers.
+The layers themselves are circular, flat, and concentrically stacked, forming an $l$-stepped pyramid with a flat top.
+While atoms within each layer occupy a single plane, the edge atoms are displaced downward, effectively smoothing the stepped profile of the pyramid.
+This pyramidal cluster is encapsulated by a locally stretched graphene sheet.
+The sheet’s valence bonds elongate only directly above the group of confined atoms; outside this coverage zone, the membrane remains undeformed and tightly adhered to the substrate.
+The most significant elongation, reaching up to one percent, occurs directly above the layer edges.
+The maximum number of possible layers $l_{m}$ increases monotonically with the total number of encapsulated atoms, $N_{g}$.
+For argon atoms, only single-layer stationary states ($l_{m}=1$) are possible for $N_{g}<145$; for $145<N_{g}<550$, both single-layer and double-layer states exist ($l_{m}=2$); for $550<N_{g}<1600$, triple-layer states are added ($l_{m}=3$); for $1600<N_{g}<2500$, four-layer states appear ($l_{m}=4$); and five-layer states ($l_{m}=5$) become accessible for $2500<N_{g}<4000$.
+The graphene membrane, through van der Waals interaction with the substrate, exerts pressure on the encapsulated atomic cluster, confining it against the substrate with pressures on the order of $P\sim 1$ GPa.
+This pressure is most pronounced for single-layer atomic packings and systematically decreases as the number of layers increases.
+Thermal fluctuation modeling reveals that among all $l$-layer states of a graphene nanobubble, one ”ground” state always exist.
+This state smoothly transitions into a layerless liquid configuration as the temperature increases, whereas the remaining $l$-layer states undergo a transition into this ground state at a distinct temperature $T_{l}$.
+For a cluster of $N_{g}=4000$ encapsulated atoms, the ground state corresponds to the four-layer packing ($l=4$).
+The thermal stability of $l$-layer states varies significantly with the atomic species.
+For helium atoms, the $l$-layer states persist up to temperatures $T_{l}=290$, $72$, $47$, and $23$ K for $l=1$, 22, 33, and $5$, respectively.
+For neon, the corresponding stability limits are 390, 120, 90, and 60 K; for argon—630, 270, 180, and 120 K; for krypton—720, 300, 240, and 180 K; and for xenon—780, 390, 330, and 240 K.
+The coexistence of multiple stable states with distinct layer numbers at low temperatures demonstrates that universal scaling does not apply to these graphene nanobubbles.
+Specifically, the height-to-radius ratio $H/R$ is not constant and can vary across the range from 0 to 0.28.
+The commonly reported universal relation $H/R\approx 0.2$ holds only for the ground states of the system.
 Acknowledgements
-The research was funded by the Russian Science Foundation (RSF) (project No. 25-73-20038).
+Computational facilities were provided by the Joint Supercomputer center (JSCC) of the National Research Center ”Kurchatov Institute”.
+The research was funded by the Russian Science Foundation (RSF) (project No.
+25-73-20038).
+Data Availability
+All data are openly available data